A Bibliography of Publications in Biometrics: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

26 December 2021 Version 1.01

Title word cross-reference

2^{M} [450]. 2 × 2 [203, 304, 58, 66, 102]. 2 × K [235, 264]. F [13, 385, 202]. F 2 [145]. F_{2} [214]. H_{0} [230]. IIB [144]. k(2 × 2) [310]. k × 2 [312]. λ [106]. M_{t} [320]. N [77]. P [420, 222, 16]. Pr(X < Y) − Pr(X > Y) [170]. R^{2} [106]. t [13, 202]. T^{2} [106]. x^{2} [312].

-estimation [77]. -value [222, 16].

1 [325]. 10.1111/j.0006 [471]. 10.1111/j.0006-341X.1999.00603.x [471]. 1994 [175]. 1997. [470]. 1999 [469, 471, 472].

20th [175].

341X.1999.00603.x [471].

44 [49, 57]. 45 [41]. 46 [56, 88]. 47 [138].

50 [418]. 53 [470]. 55 [469, 471, 472].

89d [57]. 89h [49].

90k [41]. 92f [138].

Aalen [260]. Abrupt [420]. absolute [470, 198, 221]. abundance [180]. Accelerated [385, 334, 262]. accept [230]. Access [381]. Accounting [9, 219]. accuracy [191]. Acknowledgements [468]. action [173]. Active [335]. Adaptive [294, 459, 330]. addition [173]. additional [217]. Additive [427, 398, 194, 260, 9]. additivity [226]. adequacy [265]. Adjust [279]. Adjusted [331, 428, 341]. Adjustment [407, 234]. aerial [168]. Affected [386, 266]. after [273]. against [155]. Age [432, 260, 386]. Age-At-Onset [432, 386]. Age-dependent [260]. aggregate [187]. agreement [15, 253, 122]. agricultural [46]. AIDS [116, 394, 399, 77, 325]. Albert [471]. Algorithm [332, 112, 109, 215, 201]. Algorithms [346]. alignments [159]. All-Way [450]. Allocations [464]. Allometric [443]. Allowing [404, 218]. Alternative [306, 361, 222]. Alternatives [374]. analyses [258, 117]. Analysing [146, 169]. Analysis [284, 333, 353, 290, 372, 303, 12, 357, 439, 239, 128, 278, 424, 445, 358, 422, 335, 349, 140, 396, 399, 285, 454, 330, 156, 223, 171, 68, 209, 210, 190, 7, 247, 82, 79, 85, 107, 212, 148, 203, 114, 26, 152, 35, 48, 251, 191, 51, 196, 153, 167, 243, 58, 98, 266, 122, 90, 172, 102]. Analyzing [310, 319, 202]. Angioplasty [333]. Angular [379, 84]. animal [180]. Anonymized [289]. ANOVA [13, 113, 285]. Anticipation [386]. Applicable [325]. Application [406, 462, 288, 356, 394, 278, 340, 347, 450, 376, 137]. Applications [342, 399, 163, 99, 268, 234, 77]. applied [89]. Applying [329]. Approach [326, 373, 394, 390, 415, 146, 114, 55, 78, 255]. Approaches [333, 433]. Approximate [367, 31, 56, 88, 299, 162]. Approximation [430]. Approximations [293, 230, 207]. aquatic [78]. Arbitrarily [404]. Arbitrary [336]. areas [252]. ARMA [183]. armed [143]. Arthur [129]. ascertained [77]. Ascertainment [432]. Assay [444, 283]. Assessing [82, 396]. Assessment [472, 173, 265, 370]. assignment [70]. associated [52]. Association [414, 280, 52, 182, 166]. Asymptotic [161, 20]. augmented [187]. authors [466]. autocorrelation [206].

B [470, 469, 463]. balanced [94, 85]. Bands [427, 161]. Bandwidth [377]. bandwidths [141]. Based [442, 453, 317, 433, 306, 418, 446, 374, 398, 121, 230, 146, 148, 103, 250, 270, 244, 237, 241]. Baseline [382]. Baselines [311]. Bayes [419, 11, 417]. Bayesian [471, 463, 171, 371, 151, 434, 269, 352, 326, 89, 373, 180, 304, 424, 271, 390, 132, 335, 349, 250, 144, 172, 334]. behavior [222]. Between [369, 218, 173, 52, 27, 244, 166]. Bias [432, 345, 283, 124, 168, 252]. Bias-Corrected [283]. biased [23, 55]. Binary [469, 329, 420, 366, 295, 347, 376, 286, 448, 302, 280, 451, 49, 147, 18, 146, 169, 99, 130, 93, 205, 236]. Binomial [442, 31, 56, 88, 60, 32, 59, 40, 38]. bioassay [6, 80]. Bioassays [419]. Biodiversity [462]. Bioequivalence [464, 113]. Biological [353]. Biometrics [138, 49, 56, 88, 41, 57, 465, 412]. Biometry [315]. Biostatistical [3]. Bivariate [281, 280, 63, 117, 182, 220]. blocks [179]. Blood [357]. Bone [273]. Book [314, 363, 413, 467]. Bootstrap [138, 65]. Borgan [470]. both [117]. bound [10]. boundaries [73]. Bounds [96, 423]. Breeding [395]. Bus [381]. Bus-Route [381].

C [469, 128]. Calculations [459, 72]. Calibration [241, 225]. Cambridge [42]. Cancer [378, 4]. Capture [371, 309, 322, 450, 320, 8, 30, 248, 189, 196, 69]. capture-mark-recapture [189]. Capture-recapture [8, 30, 248, 196, 69]. Carcinogenicity [456]. Carlo [350, 299, 321, 271, 172]. Case [472, 273, 432, 390, 370, 470, 223, 50, 200, 26, 221, 51, 62]. Case-Control [472, 390, 370, 470, 200, 26, 221, 51, 62]. Case-Sibling [432]. Cases [461]. catchabilities [160]. Categorical [404, 295, 393, 336, 157, 261]. categories [235, 126]. Category [404]. Causal [157, 70]. Cells [342]. Censored [471, 327, 298, 317, 352, 329, 356, 394, 337, 426, 447, 338, 349, 399, 341, 199, 165, 82, 269, 192, 268, 39, 177, 167, 262]. Censoring [339, 141, 1, 2, 41, 57, 237]. Censorship [429]. census [238, 12]. Censuses [313]. Chain [321, 271, 447, 339, 164, 172]. Chain-Of-Events [447, 339]. chains [205]. Change [329]. Changepoint [420]. changes [1, 2, 41, 57]. character [135]. chi [74]. chi-square [74]. Child [386]. Choice [461]. Choices [404]. Chromosome [277]. Circular [379, 96, 123]. Clarice [472]. Class [436, 107, 191]. Classical [89, 407]. classification [47, 229]. classified [142]. Clinical [291, 389, 435, 331, 292, 335, 409, 440, 438, 325, 163, 119, 181, 79, 143, 76, 117, 5, 4, 250, 98, 144]. Closed [440, 238, 83, 248]. Closure [320]. Cluster [310, 333, 231, 170]. Clustered [333, 375, 347, 376, 383]. clustering [121, 20]. Coal [361]. Coefficient [455, 154]. Coefficients [405, 258, 10, 166]. Cohen [15]. cohort [156, 120, 97]. collaboration [3]. combination [101]. Combinations [419]. Combined [154, 173, 16]. Combining [304, 390]. comes [376]. Comment [464, 138]. Commonly [303]. Comparing [288, 426, 252, 165, 246]. Comparison [333, 22, 462, 251, 438, 359, 50, 21, 253, 66, 1, 2, 41, 57]. Comparisons [316, 374, 230, 185, 143]. Competition [305]. completely [142]. Complex [365]. Complications [333]. Component [354, 190, 127, 16]. Components [284, 290, 112, 184, 100, 10]. computing [12]. Concentration [283, 173]. Conception [415]. Condition [381]. Conditional [299, 304, 405, 230, 12, 81, 2, 41, 102]. Conditions [287]. Confidence [351, 327, 442, 460, 24, 427, 305, 283, 242, 185, 214, 161, 207, 208, 99, 91, 10, 170]. confounders [81]. Confounding [333, 233]. consensus [16]. considerations [181]. constant [15, 137]. constrained [124]. constructing [164]. Construction [214]. Contain [359]. Contingency [303, 450, 209, 107, 126, 66, 142, 102]. Continual [4]. Continuous [376, 398, 192, 67, 237]. Continuous-Time [398]. continuously [52]. Contoured [457]. contribution [106]. Contributions [318, 44, 46, 43]. Control [472, 414, 432, 390, 388, 335, 456, 370, 470, 200, 26, 221, 246, 80, 51, 62]. controls [68, 13]. controversy [44]. Convergence [194]. Cook [124]. copula [182]. Coronary [333]. correct [168]. Corrected [283]. Correction [470, 469, 471, 472, 49, 56, 88, 288, 41, 57, 31]. Correlated [379, 448, 302, 224, 186, 90, 236, 252]. Correlation [469, 286, 369, 401, 22, 193]. count [239, 168, 33]. counts [243]. Covariance [439, 407, 111, 194, 206, 33, 9]. covariances [67]. Covariate [385, 329, 336, 267, 120]. Covariates [324, 350, 295, 410, 383, 341, 92, 261, 86, 52]. coverage [160]. Cox [245, 430, 329, 267, 294, 261, 340, 348, 270]. criterion [48]. critical [119, 70]. cross [47, 150, 142]. cross-classification [47]. cross-classified [142]. cross-sectional [150]. crosses [179]. crossover [94, 35]. Cumulative [427, 249, 211, 126]. cured [177]. Curve [407, 171, 53, 166]. Curves [290, 427, 401, 330, 37, 64, 252]. Cycles [403].

D [469, 471, 128]. Daily [403, 415]. Data [452, 469, 471, 327, 333, 298, 291, 437, 371, 385, 382, 317, 346, 352, 288, 360, 439, 366, 272, 322, 356, 394, 337, 444, 458, 425, 445, 410, 316, 426, 344, 388, 431, 347, 403, 422, 286, 383, 393, 447, 338, 436, 448, 349, 339, 399, 456, 302, 280, 336, 411, 397, 454, 325, 279, 311, 341, 199, 470, 163, 49, 147, 18, 61, 7, 131, 22, 82, 269, 192, 116, 258, 162, 85, 227, 169, 259, 158, 63, 256, 239, 268, 39, 221, 164, 34, 120, 177, 35, 251, 73, 80, 182, 220, 263, 93, 167, 186, 33]. data [90, 172, 134, 97, 142, 228, 237, 236]. Database [364]. Dates [431]. David [472, 139]. Decision [409, 250]. defined [135]. Degree [425]. Delay [317]. Deleterious [321]. delivered [175]. delta [36]. Demétrio [469]. Densities [377, 207, 24]. Density [152]. Dependence [425]. Dependent [429, 318, 49, 260]. describing [29]. Design [441, 418, 335, 323, 181, 185, 37, 152, 4, 45, 58, 98]. designed [7]. Designing [381, 292]. Designs [218, 402, 392, 409, 328, 199, 94, 204, 103, 35, 6, 93, 38]. desirable [101]. detailed [234]. Detecting [64, 145, 127]. determination [120, 250]. Determining [73]. development [29]. Developmental [376]. deviation [198]. Diagnosing [212]. Diagnostic [448, 191, 252]. Diagnostics [430]. Diagonals [126]. Diagonals-parameter [126]. diallel [179]. dichotomous [17]. Dietary [357]. Difference [442, 31, 56, 88, 110]. Differences [319]. Diggle [288]. Dilated [345]. Dilution [444, 283, 133]. Dimensional [374, 210, 142]. Dinosaur [462]. directed [264]. Dirichlet [219]. Discrete [387, 383, 339, 256, 67, 263, 186, 69]. Discrete-Time [339]. Discriminant [290, 223, 212, 48]. Disease [353, 433, 441, 20]. Diseases [293, 432]. Disequilibrium [455]. dispersion [60, 40]. dissimilarity [244]. distance [107, 178, 244]. Distances [423, 27]. Distribution [317, 229, 428, 341, 32, 219, 20, 232, 174]. Distribution-free [229]. Distributions [457, 396, 167]. Diversity [318]. Diversity-Dependent [318]. DNA [149, 423, 364, 27, 244]. DOI [471]. Dose [449, 453, 101, 37, 80]. Dose-Response [453, 37]. Double [365, 168]. double-count [168]. Double-Platform [365]. doubling [240]. Doubly [399, 209]. Drug [396]. Dual [391, 94]. dual-balanced [94]. Due [333]. d’une [95]. Dunnett [50]. Duration [463, 5]. Durations [301, 300]. Dynamics [325, 29].

Early [230, 438, 247, 5]. Ecological [288]. Ed [418]. Edge [288]. Edge-Effect [288]. Editor [465, 412]. Effect [288, 273, 259, 113]. Effective [449]. Effects [333, 291, 372, 356, 337, 407, 378, 345, 305, 282, 417, 280, 285, 279, 224, 49, 18, 173, 162, 47, 160, 193, 233, 34, 191, 81, 70, 188, 186, 75, 201, 9]. Efficacy [380, 374, 117]. Efficiency [472, 380, 393, 370, 35, 87]. Efficient [37, 428, 341]. Effort [381]. Eggs [287]. either [165]. Elevations [373]. Elliptically [457]. emphasis [44]. Empirical [333, 419, 417]. ended [235]. Endpoints [440, 438, 117, 202, 98, 255]. Entropies [462]. Environment [431, 247]. epidemic [95]. epidémie [95]. epidemiology [234]. Epistatic [319]. equality [206]. equation [69]. Equations [387, 336, 411, 397, 454, 261, 67]. equicorrelation [154]. Equivalence [445, 19, 99, 125]. Error [276, 324, 354, 423, 448, 226, 207, 267, 26, 234, 86, 228, 87]. Errors [367, 395, 257]. Estimate [381, 378, 80, 178]. Estimated [396, 208, 170]. Estimates [446, 158, 268, 154]. Estimating [469, 353, 380, 433, 259, 187, 441, 113, 267, 261, 425, 316, 67, 387, 286, 81, 296, 447, 336, 411, 397, 454, 69]. Estimation [245, 470, 313, 298, 419, 375, 317, 346, 104, 377, 322, 297, 221, 432, 418, 368, 450, 150, 339, 282, 417, 159, 369, 1, 2, 41, 57, 398, 428, 341, 15, 112, 95, 224, 83, 111, 85, 31, 56, 88, 71, 91, 37, 180, 193, 152, 59, 141, 40, 248, 178, 215, 110, 77, 8]. Estimator [416, 39]. estimators [22, 160, 123, 66, 262]. Evaluating [253, 364, 191]. Evaluation [233, 234]. Event [471, 290, 352, 243]. Event-Related [290]. Events [447, 339, 77]. Evidence [364]. Evolution [315, 149]. Exact [242, 442, 184, 47, 208, 202, 355, 100, 450, 305, 401, 111, 203, 205, 58, 102]. Excess [451]. exchangeable [169]. existence [101]. expectation [81]. expected [13]. experiment [69]. Experimental [418, 422, 328, 45]. Experiments [398, 7, 72, 78, 196, 140]. explanatory [147]. Exploratory [128]. Exponential [299, 460, 11, 120]. Exposure [472, 281, 441, 316, 301, 370, 300, 81]. Exposure-Disease [441]. Extension [443, 50, 31, 56, 88, 29]. Extensions [316]. Extent [447]. extra [59]. extra-binomial [59]. extra-Poisson [59].

F [128]. factorial [199, 140]. Failure [291, 385, 382, 273, 425, 399, 334, 61, 263, 110, 186, 134, 225]. fairly [42]. familial [22, 258]. Families [299, 460, 193]. Family [392, 16]. family-wide [16]. fiducial [175]. Fieller [36]. Finding [453]. Finite [332, 54, 122]. Firth [469]. Fish [287]. Fisher [175, 42, 44, 45, 46, 43]. Fishing [381]. Fit [355, 348, 320, 270, 262]. Fitting [385, 431, 338, 62, 216]. Fixed [107, 47, 188]. Fixed-distance [107]. Flexible [276]. Florence [139]. Follmann [471]. follows [466]. Forage [395]. Forest [373]. Form [307]. Found [364]. Frailties [358, 245]. free [229]. French [95]. frequencies [244]. full [71, 169]. Function [298, 402, 368, 249]. Functions [293, 104, 21, 260, 125]. fundamental [175]. Fundamentals [272, 128].

G [469, 138]. gamma [245]. Gaussian [121, 194]. Gene [326, 321, 145]. General [284, 156, 53, 196]. Generalized [284, 354, 336, 161, 162, 219, 17, 155, 215]. generation [247]. genes [240]. Genetic [277, 278, 423, 176, 178, 9]. genetical [43]. geneticist [42]. genetics [42, 44]. Genomic [414]. Genotype [431, 28]. Genotype-By-Environment [431]. Genotypes [318, 240]. genotypic [135]. Geographic [378]. geometric [152]. Gibbs [95]. Global [130, 348]. gold [253]. Goodness [355, 348, 320, 270]. Goodness-of-Fit [348, 320, 270]. Graft [273]. Grasses [431]. Gregor [44]. Group [389, 297, 117, 392, 459, 306, 204, 440, 79, 208, 72]. Grouped [383, 61, 192, 63, 120, 38, 134]. groups [50]. Growth [407, 397, 401, 330, 171, 7, 53, 64, 189, 166]. guidelines [138, 65, 144].

Haemophilus [463]. Haenszel [211, 134, 236]. Haenszel-type [211]. Hall [138]. haphazard [187]. Hardy [455]. Hardy-Weinberg [455]. Hatching [287]. Hazard [298, 293, 141, 274, 87]. Hazards [437, 382, 329, 398, 249, 116, 71, 133, 21, 137, 86, 197, 115, 270]. Heading [431]. Healing [406]. health [137]. heredity [44]. Heritabilities [395]. heritability [242]. Heterogeneity [309, 344, 76]. heterogeneous [231]. Heteroscedastic [285]. Hidden [345]. Hierarchical [463, 437, 421, 304, 390, 11]. High [210]. High-dimensional [210]. Highly [437, 384]. Historical [456, 68, 80]. History [175]. HIV [281, 289, 356, 150, 325]. HIV-1 [325]. Hoaglin [128]. Homogeneity [308, 147, 176]. homoscedasticity [229]. Hotelling [106]. Household [380]. hypotheses [32, 155]. Hypothesis [278, 292, 138, 65, 222, 197].

Identification [449, 28]. Identifying [289]. II [181, 218, 250, 409]. immune [177]. Immunity [463]. immunoassay [241]. Impact [357, 76]. implications [70]. Improve [472, 370]. Improved [423]. Improvements [256]. Improving [109]. Incidence [353, 433, 446, 427, 249, 150]. including [186]. Inclusion [289]. Incomplete [385, 439, 295, 410, 146, 261, 179]. Incorporating [181, 68]. increase [13]. independence [220, 102]. Independent [375, 274, 359, 173]. index [54, 20]. Indicators [387].

Individual [455, 380, 445, 251, 106]. Individuals [289, 177]. Induced [429]. Infant [330]. Infection [289, 150]. Infectiousness [380]. Inference [371, 434, 372, 299, 444, 271, 292, 401, 166, 199, 175, 211, 132, 108, 70, 153, 241]. Inferences [182]. Inferential [325]. inflation [174]. influence [82]. Influential [461, 212, 64]. Influenzas [463]. information [217]. Informative [279, 162, 266, 1, 2, 41, 57]. informatively [167]. Initiation [338]. Intake [357, 369]. Interaction [435, 343, 450]. Interactions [421, 105]. intercept [136]. interclass [22]. interim [117]. Interlaboratory [285]. Intermittent [346]. Interpretation [258]. Interspecific [319]. Interval [298, 83, 329, 450, 447, 338, 305, 349, 339, 192, 31, 56, 88, 243, 10, 237, 170]. Interval-Censored [298, 447, 338, 349]. Intervals [351, 327, 442, 283, 242, 185, 214, 207, 208, 99, 91, 24]. Intraclass [469, 286, 401, 193]. invariant [192]. invariants [108]. inverse [9]. Investigation [94]. ion [387]. irreducible [164]. Isotonic [406]. Issues [357].

J [128]. j.0006-341X.1999.00603.x [471]. Jacobson [124]. joint [78, 228]. Jolly [160]. June [470, 471].

kappa [15]. Kernel [377, 254]. kernels [141]. Key [336]. Known [416, 42, 8]. Knox [343]. Kriging [368].

lack [262]. lack-of-fit [262]. Langholz [470]. Latent [408, 387, 188, 448, 302, 107, 191, 122]. layouts [246]. least [198, 35]. Lecture [175]. Left [317, 165]. left- [165]. Left-Censored [317]. length [163]. Leslie [29]. lethal [37]. Letter [412]. Leukemia [273]. Levels [356]. Life [426, 262]. Lifetime [458, 341, 104]. Likelihood [284, 394, 295, 418, 376, 365, 364, 285, 15, 224, 36, 146, 85, 71, 169, 127, 108, 40, 154]. Likelihood-Based [418, 146]. Likelihoods [307, 304]. Limited [313]. Limiting [193]. Linder [129]. Line [375, 377, 365, 274, 257, 180, 254, 136, 24, 55]. Linear [284, 231, 354, 355, 223, 242, 224, 162, 235, 89, 268, 137, 234, 108, 100, 17, 155, 232, 2, 41]. Link [237]. Link-based [237]. Linkage [278, 176, 266]. Literature [445]. Lives [396]. Local [298, 71, 377, 268]. Locally [341]. location [214]. loci [214]. locus [28]. Log [355, 232, 359, 385, 125, 87]. log- [385]. Log-linear [232]. log-rank [125, 87]. Logistic [30, 366, 417, 328, 111, 258, 207, 62]. Logistic-Normal [366]. Logit [351, 309, 52]. logit-rank [52]. London [42, 175]. Longitudinal [452, 434, 406, 366, 295, 280, 146, 85, 259, 132, 33, 228]. lower [10].

M [469, 472]. Mahalanobis [244]. major [28, 145]. Mann [25]. Mantel [211, 134, 236]. Many [404]. March [469]. Marginal [382, 406, 207]. Marginally [366]. Margins [360]. mark [83, 158, 189]. mark-recapture [83, 158]. Marker [452, 446, 150]. Marker-Based [446]. Markers [326, 266]. Markov [400, 346, 321, 265, 164, 271, 338, 205, 339, 172]. Markovian [149].

Marrow [273]. Mars [424]. Martingale [322, 270, 69]. matched [200, 26, 51, 58]. matched-pairs [58]. Matching [384]. matrices [206]. Matrix [316, 29, 9]. Maximally [312, 74]. Maximum [15, 284, 224, 40, 285, 85, 108, 154]. McNemar [205]. Mean [402, 104, 26]. Means [359, 113, 67]. measure [244]. Measured [324, 267, 52, 228]. Measurement [276, 257, 354, 226, 207, 234, 86]. Measurements [379, 457, 422, 157, 64]. Measures [369, 279, 131, 114, 34, 72, 115, 90, 75]. mechanism [70]. median [37, 110]. medical [3]. member [193]. Memorial [175]. Memory [361]. Mendel [44]. Meta [372, 445, 68, 251]. Meta-Analysis [372, 445, 68, 251]. Metaregression [357]. Method [379, 289, 343, 282, 156, 36, 269, 136, 4, 253, 195]. methodology [196]. Methods [281, 322, 295, 271, 316, 426, 431, 365, 227, 234, 254, 59, 168]. Minimally [420]. Minimum [449]. Misclassified [441, 316]. Mismeasurement [281]. Missing [346, 350, 301, 336, 411, 162, 300]. Missingness [336, 279]. misspecification [60, 87]. Mixed [382, 400, 307, 309, 356, 354, 407, 378, 431, 282, 417, 280, 75, 242, 224, 111, 47, 89, 113, 231, 34, 100, 167, 215, 186]. mixed-effect [113]. Mixed-effects [75]. Mixed-Form [307]. Mixture [278, 410, 332, 122]. Mixtures [318, 239, 127]. Model [452, 463, 121, 291, 430, 385, 382, 400, 329, 439, 394, 337, 294, 304, 424, 378, 431, 387, 361, 275, 319, 427, 448, 349, 320, 302, 282, 334, 461, 15, 245, 171, 68, 242, 249, 49, 18, 7, 111, 124, 192, 116, 162, 47, 207, 71, 89, 113, 267, 133, 149, 53, 265, 261, 64, 120, 260, 86, 197, 167, 29, 186, 115, 126, 122, 270, 2, 41, 228, 262, 87, 166]. Model-based [121]. Modeling [404, 408, 373, 337, 332, 344, 403, 383, 415, 263, 311, 95, 232, 1, 57]. Modélisation [95]. Modelling [85, 362, 81]. Models [284, 437, 276, 309, 277, 346, 408, 367, 366, 356, 350, 273, 278, 354, 458, 355, 410, 407, 348, 405, 431, 376, 338, 274, 339, 417, 280, 397, 142, 325, 398, 328, 216, 30, 163, 112, 224, 157, 161, 226, 183, 84, 107, 21, 265, 231, 34, 194, 100, 17, 191, 188, 62, 182, 155, 215, 33, 201, 97, 237]. Modern [407]. modes [70]. Modification [389]. modifications [256]. modified [215]. Modifying [13]. molecular [159]. moment [59]. Monitoring [471, 406, 143, 352, 19, 204]. monotone [115]. Monte [350, 299, 321, 271, 172]. Mortality [433]. Mosteller [128]. Motility [342]. Motion [342]. Mover [452]. MR0931633 [57]. MR0963914 [49]. MR1029611 [41]. MR1132543 [138]. multi [247, 143]. multi-armed [143]. multi-environment [247]. Multicenter [291, 331]. multidimensional [259]. multifactor [105]. Multinomial [328, 61, 232]. Multiple [163, 462, 281, 304, 387, 409, 440, 118, 50, 185, 31, 56, 88, 12, 256, 202, 130, 98, 195, 241]. multiple-recapture [12]. multiple-step [118]. Multiplicative [319, 234]. Multiplicity [408]. Multipoint [266]. Multistage [353, 54, 119]. multistratum [90]. Multitype [421]. Multivariate [291, 419, 382, 435, 337, 457, 425, 407, 358, 330, 163, 210, 112, 73, 67, 78, 263, 186, 166]. Museum [175]. mutation [9].

natural [175, 175]. negative [60, 40, 232]. Nest [344]. Nested [273, 470, 221]. Nightingale [139]. no [138, 49, 56, 88, 41, 57]. non [121]. non-Gaussian [121]. Nonignorably [411]. Nonlinear [34, 431, 161, 75, 201]. Nonmonotonic [453]. Nonparametric [199, 209, 449, 317, 373, 268, 52, 368, 123, 339, 110, 438, 216, 165, 91, 55, 255]. nonresponse [157]. Nonzero [311]. Normal [430, 366, 337, 359, 99, 127, 167]. Note [305, 203, 136]. Number [313, 416]. Nurseries [395]. Nutrient [369].

Obituary [129, 14, 139]. Object [384]. Observation [400]. Observational [345]. Observations [359, 165, 212]. Observer [375]. Observers [274]. occupational [137, 234]. October [175]. Odds [351, 441, 316, 211, 66]. off [218]. Offspring [313]. omitting [48]. Oncology [374]. One [17, 292, 285, 246, 229, 155]. One-Sided [292, 17, 155]. One-Way [285, 246, 229]. Onset [408, 432, 386]. Open [322, 235, 196]. open-ended [235]. Optimal [402, 464, 38, 318, 151]. Optimization [119]. Optimum [328]. order [264, 217]. order-directed [264]. Ordered [393, 209, 104, 235, 264, 126]. Ordering [326]. Ordinal [336, 311, 211, 186, 115]. ordinary [35]. Other [420, 343]. Outcome [281]. Outcomes [332, 409, 130, 195]. Outcrossing [319]. Outlier [340]. overdispersed [239]. Overdispersion [287, 33]. Overlap [305]. overviews [76].

P [471, 138]. pair [26]. pair-matched [26]. paired [205]. Pairs [386, 58, 266]. pairwise [185, 143]. Panel [400]. Parameter [283, 158, 60, 40, 182, 126]. Parameters [425, 417, 280, 397, 31, 56, 88, 267, 67]. Parametric [276, 350, 216, 39]. Parent [386]. Partial [289, 179]. partially [39, 266, 142]. Partly [405]. Patient [445, 251]. Pattern [410]. Pattern-Mixture [410]. Patterns [421, 336]. pedigree [164]. Penalized [394]. Penetrance [277]. percent [87]. Percentile [92]. Performances [288]. Permutation [198, 379, 436]. Permutation-Type [379]. permutational [195]. Petersen [416]. pharmaceutical [151]. Phase [434, 181, 4, 144, 218, 250, 409]. phenotypic [135]. Phylogenetic [271, 108]. Piecewise [21, 92]. pilot [103]. Planning [19]. Plants [321]. Platform [365]. plot [123]. Point [421, 254, 66]. Poisson [373, 239, 300, 59, 301, 174]. polygenic [194]. Polyhazard [458]. polynomials [92]. Pooled [472, 158, 370]. Pooling [353]. Population [433, 377, 322, 450, 296, 325, 359, 398, 163, 158, 145, 248, 189, 29]. Population-Based [433]. Populations [375, 446, 238, 214, 83, 54, 196]. portmanteau [183]. Positive [358]. Possibly [359]. Potential [290]. Power [200, 308, 230, 131, 134]. powerful [127, 25]. Practical [70, 144, 4]. Predict [463]. predicting [11]. Prediction [249, 405, 89]. Predictive [360, 15, 132]. preferred [118]. Presence [438, 213, 39, 177, 8, 1, 2, 41, 57, 252]. Pressure [357]. Prevalence [353, 434, 433]. prevalent [97]. Prevosti [178]. Principal [290, 190]. Probabilities [327, 415, 12, 187, 96, 204, 126]. probability [15, 23]. Probit [324, 210, 224]. probit-linear [224]. Problem [429, 426, 137]. procedure [118, 50, 151, 79, 169, 256]. Procedures [306, 440, 138, 11, 52]. Process [406, 95, 1, 57]. Processes [400, 239]. processus [95]. Production [313]. professor [42]. progeny [28]. progeny-tested [28]. Progressive [293, 394]. Propelled [342]. Properties [438]. Proportional [437, 329, 86, 115, 249, 116, 71, 133, 21, 197, 270]. Proportions [442, 297, 208]. Prospective [390, 308]. protection [118].

quadratic [212]. Qualitative [105]. Quality [388, 428, 341]. Quality-Adjusted [428, 341]. Quanta1 [323]. quantal [37, 80]. Quantifying [160]. Quantiles [327, 345]. quantitative [214].

R [472, 138]. R. [42, 45, 43]. Random [291, 372, 337, 191, 417, 285, 279, 163, 224, 49, 18, 162, 47, 141, 201]. Randomization [380]. Randomized [357, 98]. randomly [39]. Randomness [288]. Rank [453, 192, 306, 52, 78, 125, 87]. Rank-Based [453, 306]. Ranked [23]. Ranking [289]. Rat [403]. ratchet [96]. Rate [432, 141, 189]. Rates [471, 352, 355, 378]. rating [122]. Ratio [351, 441, 432, 242, 135, 127, 110, 66, 8, 87]. ratio-defined [135]. Ratios [316, 364, 211]. real [73]. reassessment [4]. Recapture [371, 309, 322, 450, 320, 397, 398, 30, 83, 12, 158, 248, 189, 196, 8, 69]. recaptures [187]. recovery [238, 172]. Recreational [381]. recruitment [189]. recurrent [243]. Reduced [277, 345, 78]. reduced-rank [78]. Reducing [26]. Reflect [309]. Regions [460]. Registries [433]. Regression [437, 384, 406, 84, 324, 63, 133, 350, 135, 337, 294, 425, 340, 348, 405, 431, 347, 331, 301, 399, 417, 225, 311, 451, 216, 30, 245, 198, 161, 82, 258, 207, 268, 265, 300, 234, 197, 78, 62, 243, 75]. regressive [258]. Related [277, 290, 190]. Relating [407]. relation [15]. Relationship [357, 173]. Relationships [441]. Relative [308, 13, 266]. reliability [159]. REML [112, 111]. Removal [454]. Repeated [324, 289, 457, 422, 279, 157, 131, 114, 99, 34, 72, 115, 90, 75]. repeated-measures [72]. Repeatedly [446]. replied [466]. reply [138]. Reports [289]. research [3]. residual [104]. Residuals [340, 270]. Resistant [384]. Response [453, 287, 323, 280, 411, 218, 84, 117, 37, 5, 211, 17, 80, 186, 115]. Responses [295, 347, 336, 146, 99, 67, 186]. Restricted [369]. results [76, 246]. retrospectively [77]. reuse [227]. Review [128, 45]. Reviews [314, 363, 413, 467]. Ridout [469]. Right [317, 341, 165, 1, 2, 41, 57]. Right-Censored [341, 165]. Right-Truncated [317]. ring [172]. ring-recovery [172]. Risk [427, 470, 221, 260]. Risks [308, 451]. RNA [356]. Robust [59, 347, 393, 93, 114]. Robustness [60, 297, 302]. ROC [233, 252]. role [70]. Route [381]. rule [5]. rules [76]. runs [241].

S [469, 471, 138]. Saddlepoint [207, 293]. Safety [374, 117]. Sample [371, 389, 429, 460, 459, 120, 426, 308, 72, 250, 238, 61, 151, 131, 200, 227, 160, 91, 132, 240, 262]. sample-size-optimal [151]. Samples [351, 288, 359, 165, 49, 187, 170]. Sampling [310, 377, 178, 369, 27, 180, 254, 23, 136, 24, 55]. Saved [426]. scale [262]. scan [96]. scheme [164]. score [264, 174]. Screened [446]. Screening [391, 340]. Search [364]. Seber [160]. sectional [150]. Selected [312, 420, 74]. Selection [377, 103, 410, 349, 118, 175, 269, 53, 23, 54, 217]. Self [342, 292]. Self-Designing [292]. Self-Propelled [342]. Semi [338, 339]. Semi-Markov [338, 339]. Semiparametric [367, 281, 439, 243, 282, 334]. Sensitivity [51, 345]. September [472]. Sequence [420, 159, 244]. Sequences [423, 27]. Sequential [389, 99, 392, 459, 306, 6, 426, 440, 118, 151, 79, 117, 204, 38]. Serial [444]. series [265]. set [23]. sets [107]. Settings [453]. Several [333, 303, 359, 50, 203, 29, 66, 102]. Sewall [14]. sex [8]. shape [190]. shape-related [190]. shaped [260]. shared [245]. Shelf [396]. Sibling [432]. Sided [292, 17, 155]. Sign [362, 213]. significance [17, 16, 178]. SIMEX [354]. Simple [379, 11, 170]. Simulated [365]. simulation [172]. Simultaneous [185]. Single [420, 193, 66]. single-member [193]. singularities [219]. sires [28]. Site [381]. Size [190, 389, 131, 322, 459, 308, 450, 296, 398, 151, 200, 120, 23, 248, 55, 72, 240, 250]. Size- [190]. size-biased [23, 55]. skewness [31, 56, 88]. Sleep [403]. Small [351, 371, 288, 460]. Small-Sample [371, 460]. Smoothed [290]. smoothers [268]. smoothing [92]. Sodium [357]. Some [246, 33, 262, 155]. Space [343]. Spatial [247, 288, 421, 183, 227]. spatially [90]. spatio [95, 95]. Spatio-temporal [95]. spatio-temporelle [95]. special [223, 44]. Species [319, 29]. Specific [451]. Specified [366]. Specimens [353]. Sperm [342]. Spline [148]. Spline-based [148]. Splines [294, 330]. SPRT [60]. square [74, 126]. squared [26]. squares [35]. Stability [396, 153]. Stable [358]. Stage [313, 323, 181]. Stage-Structured [313]. stages [29]. Standard [367, 395, 253]. State [400, 394]. Statistic [348, 96]. Statistical [28, 357, 444, 388, 168, 153, 292, 97, 70]. Statistics [312, 74, 46, 43, 236]. Stayer [452]. step [118, 246, 195]. step-up [246, 195]. Stochastic [346, 397]. stopping [230, 76, 5]. strategies [27]. Stratified [238, 437, 248, 416, 436, 21, 211, 62]. stratum [73]. Structural [273]. Structure [321, 401]. Structured [313]. structures [194]. Studies [472, 353, 309, 414, 434, 441, 432, 390, 308, 345, 464, 323, 399, 456, 285, 461, 370, 83, 19, 200, 113, 26, 202, 233, 103, 120, 51, 105, 62, 253]. Study [273, 5, 321, 287, 189, 255]. Subdaily [403]. Subject [339, 451, 86]. Subject-Specific [451]. Subset [217]. substances [173]. Summarizing [342]. Summary [279]. Surfaces [287]. Surrogate [438, 255]. Survey [360]. Surveys [381, 375, 365, 274, 257, 132, 123, 168, 150]. Survival [327, 272, 293, 306, 424, 344, 358, 383, 393, 436, 349, 461, 428, 269, 11, 148, 5, 39, 21, 204, 103, 177, 52, 197, 182, 220, 140, 255, 228, 237]. survivor [125]. Susceptibility [380]. Suspect [364]. Symmetric [327]. symmetry [126].

Table [450, 66]. Tables [310, 312, 303, 304, 209, 107, 31, 56, 88, 235, 203, 264, 126, 142, 102]. Tag [397, 238]. Tag-Recapture [397]. tag-recovery [238]. Tags [416]. tail [96]. Tarone [256]. technique [168]. Techniques [407, 216]. temporal [95]. temporelle [95]. Test [442, 455, 147, 435, 235, 392, 306, 331, 362, 450, 305, 320, 456, 438, 165, 223, 213, 183, 256, 149, 137, 264, 127, 17, 16, 205, 25, 270, 134, 125, 236, 87, 174]. Test-Based [442]. tested [28]. Testing [379, 101, 297, 278, 177, 206, 308, 386, 440, 138, 119, 65, 208, 99, 197, 178, 195]. Tests [327, 453, 303, 288, 420, 32, 343, 53, 176, 354, 355, 393, 220, 436, 13, 198, 95, 184, 61, 131, 192, 47, 28, 148, 117, 202, 130, 246, 52, 100, 229, 191, 16, 155, 25, 262, 252]. Their [340, 359, 268]. theorem [36, 175]. Theoretic [409]. theory [178, 250]. Thermograms [275]. Three [394, 142]. three-dimensional [142]. Three-State [394]. Thresholds [451]. Thursday [175]. Ties [362, 213]. Time [291, 385, 382, 329, 273, 343, 425, 405, 383, 339, 399, 334, 398, 428, 61, 265, 52, 263, 134, 225, 69, 237]. Time-Varying [329, 405]. Times [338, 119, 110, 186]. Tits [361]. Tools [325]. Total [433]. toxicity [181, 218, 78]. Toxicology [376]. Tracking [447]. trade [218]. trade-off [218]. trait [214, 122]. Transect [375, 377, 365, 274, 257, 180, 254, 24, 55]. Transformations [226, 439]. Transient [296]. Transit [387]. Transition [408, 187]. Transplantation [273]. trapping [152]. Treatment [345, 374, 13, 259, 113]. Treatments [438, 94, 118, 246]. Trend [331, 456, 235, 137, 264]. trends [80]. Trial [388, 331, 58, 38]. Trials [291, 389, 435, 281, 357, 459, 306, 292, 335, 409, 440, 374, 438, 325, 163, 119, 181, 218, 151, 247, 79, 143, 76, 117, 5, 4, 250, 98, 144]. true [222]. Truncated [317, 394, 116]. Truncation [339, 91]. Tukey [128]. Tumor [408]. twice [42]. Two [138, 442, 61, 400, 434, 433, 65, 429, 91, 426, 323, 374, 165, 238, 209, 181, 94, 185, 131, 104, 107, 47, 127, 229, 72, 125, 262, 252]. Two-Dimensional [374]. two-group [72]. Two-Phase [434]. Two-Sample [429, 426, 61, 91, 238, 131, 262]. Two-Stage [323, 181]. Two-State [400]. two-way [209, 185, 107, 47, 229]. Type [381, 379, 211, 463]. typicality [223].

U [260]. U-shaped [260]. Umbach [472]. unbalanced [112, 185, 47, 89, 246, 100, 75]. Unconditional [303, 58]. unequal [160]. Uniformity [379]. uniformly [25]. Unifying [392]. Univariate [407, 280]. Unknown [338]. Unread [416]. Upper [345, 96]. upper-tail [96]. Use [309, 435, 456, 279, 216, 80]. Used [303]. Using [472, 290, 408, 322, 293, 273, 321, 304, 424, 445, 332, 378, 405, 370, 330, 112, 92, 152, 168, 150, 266]. Utility [307]. Utilization [189].

Vaccine [380, 281]. Valid [372]. Validating [446]. Validation [441]. value [222, 135, 16]. values [119]. variability [13]. Variable [404, 373, 432, 302, 269, 123, 188]. Variables [408, 147, 239, 267, 219, 48, 106].

Variance [284, 455, 402, 354, 368, 245, 242, 112, 184, 85, 59, 100, 251, 10, 9]. variance-covariance [9]. Variance. [128]. variances [231]. Variation [378, 59]. Variety [388, 247]. Varying [329, 405, 141]. verification [252]. Versions [213, 303]. Versus [380, 273, 251]. Via [439, 160, 271, 172]. Virological [325]. visibility [124, 168]. Vivo [325].

W [128]. Way [450, 285, 209, 185, 107, 47, 246, 229]. web [152]. Weighted [322, 32, 87]. Weinberg [472, 455]. well [42]. well-known [42]. Where [127, 165]. Whitney [25]. wide [16]. Wilcoxon [25]. Wilks [106]. Wilson [138, 138]. within [231]. within-cluster [231]. without [288]. words [244]. Wright [14].

XVIIIth [175].

Years [426].

Zero [359, 174].

References

Wu:1988:ECC

[1] Margaret C. Wu and Raymond J. Carroll. Estimation and comparison of changes in the presence of informative right censoring by modeling the censoring process. Biometrics, 44(1):175–188, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [57].

Wu:1989:ECC

[2] Margaret C. Wu and Kent R. Bailey. Estimation and comparison of changes in the presence of informative right censoring: conditional linear model. Biometrics, 45(3):939–955, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [41].

Ellenberg:1990:BCM

[3] Jonas H. Ellenberg. Biostatistical collaboration in medical research. Biometrics, 46(1):1–32, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). With discussion and a rejoinder by the author.

OQuigley:1990:CRM

[4] John O’Quigley, Margaret Pepe, and Lloyd Fisher. Continual reassessment method: a practical design for phase 1 clinical trials in cancer. Biometrics, 46(1):33–48, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:1990:SDC

[5] Kyungmann Kim and Anastasios A. Tsiatis. Study duration for clinical trials with survival response and early stopping rule. Biometrics, 46 (1):81–92, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McLeish:1990:SDB

[6] D. L. McLeish and D. Tosh. Sequential designs in bioassay. Biometrics, 46 (1):103–116, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cullis:1990:MAG

[7] B. R. Cullis and C. A. McGilchrist. A model for the analysis of growth data from designed experiments. Biometrics, 46(1):131–142, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wolter:1990:CRE

[8] Kirk M. Wolter. Capture-recapture estimation in the presence of a known sex ratio. Biometrics, 46(1):157–162, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wray:1990:AME

[9] Naomi R. Wray. Accounting for mutation effects in the additive genetic variance-covariance matrix and its inverse. Biometrics, 46(1):177–186, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1990:LBC

[10] C. Ming Wang. On the lower bound of confidence coefficients for a confidence interval on variance components. Biometrics, 46(1):187–192, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Geisser:1990:HBP

[11] Seymour Geisser. On hierarchical Bayes procedures for predicting simple exponential survival. Biometrics, 46(1):225–230, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Garthwaite:1990:AMR

[12] Paul H. Garthwaite and Stephen T. Buckland. Analysis of a multiplerecapture census by computing conditional probabilities. Biometrics, 46 (1):231–238, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brownie:1990:MAT

[13] Cavell Brownie, Dennis D. Boos, and Jacqueline Hughes-Oliver. Modifying the t and ANOVA F tests when treatment is expected to increase variability relative to controls. Biometrics, 46(1):259–266, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1990:OSW

[14] Anonymous. Obituary: Sewall Wright 1889–1988. Biometrics, 46(1): 277–279, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Aickin:1990:MLE

[15] Mikel Aickin. Maximum likelihood estimation of agreement in the constant predictive probability model, and its relation to Cohen’s kappa. Biometrics, 46(2):293–302, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rice:1990:CCV

[16] William R. Rice. A consensus combined P -value test and the family-wide significance of component tests. Biometrics, 46(2):303–308, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Piegorsch:1990:OSS

[17] Walter W. Piegorsch. One-sided significance test for generalized linear models under dichotomous response. Biometrics, 46(2):309–316, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Conaway:1990:REM

[18] Mark R. Conaway. A random effects model for binary data. Biometrics, 46 (2):317–328, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Durrleman:1990:PME

[19] Sylvain Durrleman and Richard Simon. Planning and monitoring of equivalence studies. Biometrics, 46(2):329–336, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tango:1990:ADI

[20] Toshiro Tango. Asymptotic distribution of an index for disease clustering. Biometrics, 46(2):351–357, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kronborg:1990:PCS

[21] Dorte Kronborg and Peter Aaby. Piecewise comparison of survival functions in stratified proportional hazards models. Biometrics, 46(2):375–380, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eliasziw:1990:CRE

[22] Michael Eliasziw and Allan Donner. Comparison of recent estimators of interclass correlation from familial data. Biometrics, 46(2):391–398, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muttlak:1990:RSS

[23] Hassen A. Muttlak and Lyman L. McDonald. Ranked set sampling with size-biased probability of selection. Biometrics, 46(2):435–445, 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Quang:1990:CID

[24] Pham Xuan Quang. Confidence intervals for densities in line transect sampling. Biometrics, 46(2):459–472, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Streitberg:1990:TUM

[25] Bernd Streitberg and Joachim Roehmel. On tests that are uniformly more powerful than the Wilcoxon–Mann–Whitney test. Biometrics, 46 (2):481–484, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kalish:1990:RMS

[26] Leslie A. Kalish. Reducing mean squared error in the analysis of pairmatched case-control studies. Biometrics, 46(2):493–499, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Weir:1990:SSD

[27] B. S. Weir and C. J. Basten. Sampling strategies for distances between DNA sequences. Biometrics, 46(3):551–582, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goffinet:1990:STI

[28] B. Goffinet, J. M. Elsen, and P. Le Roy. Statistical tests for identification of the genotype at a major locus of progeny-tested sires. Biometrics, 46 (3):583–594, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sondgerath:1990:ELM

[29] Dagmar Söndgerath and Otto Richter. An extension of the Leslie matrix model for describing population dynamics of species with several development stages. Biometrics, 46(3):595–607, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alho:1990:LRC

[30] Juha M. Alho. Logistic regression in capture-recapture models. Biometrics, 46(3):623–635, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gart:1990:AIE

[31] John J. Gart and Jun mo Nam. Approximate interval estimation of the difference in binomial parameters: correction for skewness and extension to multiple tables. Biometrics, 46(3):637–643, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction, [56, 88].

Kocherlakota:1990:THW

[32] S. Kocherlakota and K. Kocherlakota. Tests of hypotheses for the weighted binomial distribution. Biometrics, 46(3):645–656, 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:1990:SCM

[33] Peter F. Thall and Stephen C. Vail. Some covariance models for longitudinal count data with overdispersion. Biometrics, 46(3):657–671, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lindstrom:1990:NME

[34] Mary J. Lindstrom and Douglas M. Bates. Nonlinear mixed effects models for repeated measures data. Biometrics, 46(3):673–687, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matthews:1990:ADC

[35] J. N. S. Matthews. The analysis of data from crossover designs: the efficiency of ordinary least squares. Biometrics, 46(3):689–696, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cox:1990:FTL

[36] Christopher Cox. Fieller’s theorem, the likelihood and the delta method. Biometrics, 46(3):709–718, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kalish:1990:EDE

[37] Leslie A. Kalish. Efficient design for estimation of median lethal dose and quantal dose-response curves. Biometrics, 46(3):737–748, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Therneau:1990:ODG

[38] Terry M. Therneau, H. S. Wieand, and Myron Chang. Optimal designs for a grouped sequential binomial trial. Biometrics, 46(3):771–781, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Klein:1990:PPE

[39] John P. Klein, Shih Chang Lee, and M. L. Moeschberger. A partially parametric estimator of survival in the presence of randomly censored data. Biometrics, 46(3):795–811, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Piegorsch:1990:MLE

[40] Walter W. Piegorsch. Maximum likelihood estimation for the negative binomial dispersion parameter. Biometrics, 46(3):863–867, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:1990:CEC

[41] M. C. Wu and K. R. Bailey. Correction to: “Estimation and comparison of changes in the presence of informative right censoring: conditional linear model” [Biometrics 45 (1989), no. 3, 939–955; MR1029611 (90k:62151)]. Biometrics, 46(3):889, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [2].

Edwards:1990:RFT

[42] A. W. F. Edwards. R. A. Fisher: twice professor of genetics: London and Cambridge or “a fairly well-known geneticist”. Biometrics, 46(4):897–904, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1990:RFC

[43] E. A. Thompson. R. A. Fisher’s contributions to genetical statistics. Biometrics, 46(4):905–914, ???? 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Piegorsch:1990:FCG

[44] Walter W. Piegorsch. Fisher’s contributions to genetics and heredity, with special emphasis on the Gregor Mendel controversy. Biometrics, 46 (4):915–924, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Preece:1990:RFE

[45] D. A. Preece. R. A. Fisher and experimental design: a review. Biometrics, 46(4):925–935, ???? 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Street:1990:FCA

[46] Deborah J. Street. Fisher’s contributions to agricultural statistics. Biometrics, 46(4):937–945, ???? 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gallo:1990:ETR

[47] Jose Gallo and André I. Khuri. Exact tests for the random and fixed effects in an unbalanced mixed two-way cross-classification model. Biometrics, 46(4):1087–1095, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meulepas:1990:COV

[48] E. Meulepas. On a criterion for omitting variables in discriminant analysis. Biometrics, 46(4):1181–1183, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coffey:1990:CRE

[49] M. Coffey. Correction to: “A random effects model for binary data from dependent samples” [Biometrics 44 (1988), no. 3, 787–801; MR0963914 (89h:62166)]. Biometrics, 46(4):1247, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheung:1991:EDM

[50] Siu Hung Cheung and Burt Holland. Extension of Dunnett’s multiple comparison procedure to the case of several groups. Biometrics, 47(1):21–32, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:1991:SAM

[51] Paul R. Rosenbaum. Sensitivity analysis for matched case-control studies. Biometrics, 47(1):87–100, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OQuigley:1991:NTA

[52] John O’Quigley and Ross L. Prentice. Nonparametric tests of association between survival time and continuously measured covariates: the logit-rank and associated procedures. Biometrics, 47(1):117–127, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:1991:TMS

[53] Jack C. Lee. Tests and model selection for the general growth curve model. Biometrics, 47(1):147–159, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Norell:1991:MIS

[54] Lennart Norell, Thorvaldur Arnason, and Kristinn Hugason. Multistage index selection in finite populations. Biometrics, 47(1):205–221, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Quang:1991:NAS

[55] Pham Xuan Quang. A nonparametric approach to size-biased line transect sampling. Biometrics, 47(1):269–279, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gart:1991:CAI

[56] J. J. Gart and J. Nam. Correction to: “Approximate interval estimation of the difference in binomial parameters: correction for skewness and extension to multiple tables” [Biometrics 46 (1990), no. 3, 637–643]. Biometrics, 47(1):357, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [31].

Wu:1991:CEC

[57] M. C. Wu and R. J. Carroll. Correction to: “Estimation and comparison of changes in the presence of informative right censoring by modeling the censoring process” [Biometrics 44 (1988), no. 1, 175–188; MR0931633 (89d:62109)]. Biometrics, 47(1):357, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1].

Suissa:1991:MPT

[58] Samy Suissa and Jonathan J. Shuster. The 2 × 2 matched-pairs trial: exact unconditional design and analysis. Biometrics, 47(2):361–372, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Moore:1991:REV

[59] D. F. Moore and A. Tsiatis. Robust estimation of the variance in moment methods for extra-binomial and extra-Poisson variation. Biometrics, 47 (2):383–401, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hubbard:1991:RSN

[60] Dean J. Hubbard and O. Brian Allen. Robustness of the SPRT for a negative binomial to misspecification of the dispersion parameter. Biometrics, 47(2):419–427, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:1991:TST

[61] J. A. Cook and J. F. Lawless. Two-sample tests with multinomial or grouped failure time data. Biometrics, 47(2):445–459, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scott:1991:FLR

[62] A. J. Scott and C. J. Wild. Fitting logistic regression models in stratified case-control studies. Biometrics, 47(2):497–510, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heitjan:1991:RBG

[63] Daniel F. Heitjan. Regression with bivariate grouped data. Biometrics, 47 (2):549–562, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liski:1991:DIM

[64] Erkki P. Liski. Detecting influential measurements in a growth curves model. Biometrics, 47(2):659–668, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hall:1991:TGB

[65] Peter Hall and Susan R. Wilson. Two guidelines for bootstrap hypothesis testing. Biometrics, 47(2):757–762, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See comment [138].

Walter:1991:CSP

[66] S. D. Walter and R. J. Cook. A comparison of several point estimators of the odds ratio in a single 2 × 2 contingency table. Biometrics, 47(3): 795–811, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Prentice:1991:EEP

[67] Ross L. Prentice and Lue Ping Zhao. Estimating equations for parameters in means and covariances of multivariate discrete and continuous responses. Biometrics, 47(3):825–839, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Begg:1991:MIH

[68] Colin B. Begg and Louise Pilote. A model for incorporating historical controls into a meta-analysis. Biometrics, 47(3):899–906, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yip:1991:MEE

[69] Paul Yip. A martingale estimating equation for a capture-recapture experiment in discrete time. Biometrics, 47(3):1081–1088, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rubin:1991:PIM

[70] Donald B. Rubin. Practical implications of modes of statistical inference for causal effects and the critical role of the assignment mechanism. Biometrics, 47(4):1213–1234, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gentleman:1991:LFL

[71] R. Gentleman and J. Crowley. Local full likelihood estimation for the proportional hazards model. Biometrics, 47(4):1283–1296, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rochon:1991:SSC

[72] James Rochon. Sample size calculations for two-group repeated-measures experiments. Biometrics, 47(4):1383–1398, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pla:1991:DSB

[73] Laura Pla. Determining stratum boundaries with multivariate real data. Biometrics, 47(4):1409–1422, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Koziol:1991:MSC

[74] James A. Koziol. On maximally selected chi-square statistics. Biometrics, 47(4):1557–1561, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vonesh:1992:MEN

[75] Edward F. Vonesh and Randy L. Carter. Mixed-effects nonlinear regression for unbalanced repeated measures. Biometrics, 48(1):1–17, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hughes:1992:ISR

[76] Michael D. Hughes, Laurence S. Freedman, and Stuart J. Pocock. The impact of stopping rules on heterogeneity of results in overviews of clinical trials. Biometrics, 48(1):41–53, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1992:ERA

[77] Mei-Cheng Wang and Lai-Chu See. N-estimation from retrospectively ascertained events with applications to AIDS. Biometrics, 48(1):129–141, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ryan:1992:RRM

[78] D. A. J. Ryan, J. J. Hubert, E. M. Carter, J. B. Sprague, and J. Parrott. A reduced-rank multivariate regression approach to aquatic joint toxicity experiments. Biometrics, 48(1):155–162, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Falissard:1992:NPG

[79] Bruno Falissard and Joseph Lellouch. A new procedure for group sequential analysis in clinical trials. Biometrics, 48(2):373–388, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Prentice:1992:UHC

[80] Ross L. Prentice, Robert T. Smythe, Daniel Krewski, and Mark Mason. On the use of historical control data to estimate dose response trends in quantal bioassay. Biometrics, 48(2):459–478, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robins:1992:EEE

[81] James M. Robins, Steven D. Mark, and Whitney K. Newey. Estimating exposure effects by modelling the expectation of exposure conditional on confounders. Biometrics, 48(2):479–495, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Escobar:1992:AIR

[82] Luis A. Escobar and William Q. Meeker, Jr. Assessing influence in regression analysis with censored data. Biometrics, 48(2):507–528, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cormack:1992:IEM

[83] R. M. Cormack. Interval estimation for mark-recapture studies of closed populations. Biometrics, 48(2):567–576, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fisher:1992:RMA

[84] N. I. Fisher and A. J. Lee. Regression models for an angular response. Biometrics, 48(3):665–677, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Forcina:1992:MBL

[85] A. Forcina. Modelling balanced longitudinal data: maximum likelihood estimation and analysis of variance. Biometrics, 48(3):743–750, 1992. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nakamura:1992:PHM

[86] Tsuyoshi Nakamura. Proportional hazards model with covariates subject to measurement error. Biometrics, 48(3):829–838, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zucker:1992:EWL

[87] David M. Zucker. The efficiency of a weighted log-rank test under a percent error misspecification model for the log hazard ratio. Biometrics, 48(3):893–899, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gart:1992:CAI

[88] J. J. Gart and J. Nam. Correction to: “Approximate interval estimation of the difference in binomial parameters: correction for skewness and extension to multiple tables” [Biometrics 46 (1990), no. 3, 637–643; 1085812]. Biometrics, 48(3):979, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [31].

Harville:1992:CBP

[89] David A. Harville and Alicia L. Carriquiry. Classical and Bayesian prediction as applied to an unbalanced mixed linear model. Biometrics, 48(4): 987–1003, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Verbyla:1992:AMS

[90] A. P. Verbyla and B. R. Cullis. The analysis of multistratum and spatially correlated repeated measures data. Biometrics, 48(4):1015–1032, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:1992:TSN

[91] Richard A. Johnson, Christopher H. Morrell, and Anton Schick. Twosample nonparametric estimation and confidence intervals under truncation. Biometrics, 48(4):1043–1056, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goldstein:1992:PSU

[92] Harvey Goldstein and Huiqi Pan. Percentile smoothing using piecewise polynomials, with covariates. Biometrics, 48(4):1057–1068, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sitter:1992:RDB

[93] R. R. Sitter. Robust designs for binary data. Biometrics, 48(4):1145–1155, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carriere:1992:IDB

[94] Keumhee Chough Carrière and Gregory C. Reinsel. Investigation of dualbalanced crossover designs for two treatments. Biometrics, 48(4):1157– 1164, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chadoeuf:1992:MST

[95] J. Chadœuf, D. Nandris, J. P. Geiger, M. Nicole, and J. C. Pierrat. Modélisation spatio-temporelle d’une epidémie par un processus de Gibbs: estimation et tests. (French) [Spatio-temporal modeling of an epidemic by a Gibbs process: estimation and tests]. Biometrics, 48(4):1165–1175, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krauth:1992:BUT

[96] J. Krauth. Bounds for the upper-tail probabilities of the circular ratchet scan statistic. Biometrics, 48(4):1177–1185, 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1993:SMP

[97] Mei-Cheng Wang, Ron Brookmeyer, and Nicholas P. Jewell. Statistical models for prevalent cohort data. Biometrics, 49(1):1–11, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:1993:DAR

[98] Dei-In Tang, Nancy L. Geller, and Stuart J. Pocock. On the design and analysis of randomized clinical trials with multiple endpoints. Biometrics, 49(1):23–30, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jennison:1993:SET

[99] Christopher Jennison and Bruce W. Turnbull. Sequential equivalence testing and repeated confidence intervals, with applications to normal and binary responses. Biometrics, 49(1):31–43, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ofversten:1993:ETV

[100] J. Öfversten. Exact tests for variance components in unbalanced mixed linear models. Biometrics, 49(1):45–57, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hung:1993:TED

[101] H. M. James Hung, G. Y. H. Chi, and R. J. Lipicky. Testing for the existence of a desirable dose combination. Biometrics, 49(1):85–94, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yao:1993:EAC

[102] Qing Yao and David Tritchler. An exact analysis of conditional independence in several 2×2 contingency tables. Biometrics, 49(1):233–236, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:1993:SDP

[103] P. Y. Liu, Steve Dahlberg, and John Crowley. Selection designs for pilot studies based on survival. Biometrics, 49(2):391–398, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ebrahimi:1993:ETO

[104] Nader Ebrahimi. Estimation of two ordered mean residual lifetime functions. Biometrics, 49(2):409–417, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Russek-Cohen:1993:QIM

[105] Estelle Russek-Cohen and Richard M. Simon. Qualitative interactions in multifactor studies. Biometrics, 49(2):467–477, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rencher:1993:CIV

[106] Alvin C. Rencher. The contribution of individual variables to Hotelling’s T^{2}, Wilks’ λ, and R^{2}. Biometrics, 49(2):479–489, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Formann:1993:FDL

[107] Anton K. Formann. Fixed-distance latent class models for the analysis of sets of two-way contingency tables. Biometrics, 49(2):511–521, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Navidi:1993:PIL

[108] W. C. Navidi, G. A. Churchill, and A. von Haeseler. Phylogenetic inference: linear invariants and maximum likelihood. Biometrics, 49(2): 543–555, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rai:1993:IEA

[109] S. N. Rai and D. E. Matthews. Improving the EM algorithm. Biometrics, 49(2):587–591, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Su:1993:NED

[110] John Q. Su and L. J. Wei. Nonparametric estimation for the difference or ratio of median failure times. Biometrics, 49(2):603–607, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Drum:1993:REE

[111] Melinda L. Drum and Peter McCullagh. REML estimation with exact covariance in the logistic mixed model. Biometrics, 49(3):677–689, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Calvin:1993:REU

[112] James A. Calvin. REML estimation in unbalanced multivariate variance components models using an EM algorithm. Biometrics, 49(3):691–701, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsuan:1993:ETM

[113] Francis C. Hsuan. Estimating treatment means in a mixed-effect ANOVA model for bioequivalence studies. Biometrics, 49(3):703–713, 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huggins:1993:RAA

[114] R. M. Huggins. A robust approach to the analysis of repeated measures. Biometrics, 49(3):715–720, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thornquist:1993:PHM

[115] Mark D. Thornquist. Proportional hazards model for repeated measures with monotone ordinal response. Biometrics, 49(3):721–730, 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finkelstein:1993:PHM

[116] Dianne M. Finkelstein, Dirk F. Moore, and David A. Schoenfeld. A proportional hazards model for truncated AIDS data. Biometrics, 49(3): 731–740, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jennison:1993:GST

[117] Christopher Jennison and Bruce W. Turnbull. Group sequential tests for bivariate response: interim analyses of clinical trials with both efficacy and safety endpoints. Biometrics, 49(3):741–752, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:1993:MSS

[118] T. Timothy Chen and Richard Simon. A multiple-step selection procedure with sequential protection of preferred treatments. Biometrics, 49(3): 753–761, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brittain:1993:OMT

[119] Erica H. Brittain and Kent R. Bailey. Optimization of multistage testing times and critical values in clinical trials. Biometrics, 49(3):763–772, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lui:1993:SSD

[120] Kung-Jong Lui. Sample size determination for cohort studies under an exponential covariate model with grouped data. Biometrics, 49(3):773–778, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Banfield:1993:MBG

[121] Jeffrey D. Banfield and Adrian E. Raftery. Model-based Gaussian and non-Gaussian clustering. Biometrics, 49(3):803–821, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Uebersax:1993:LTF

[122] John S. Uebersax and William M. Grove. A latent trait finite mixture model for the analysis of rating agreement. Biometrics, 49(3):823–835, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Quang:1993:NEV

[123] Pham Xuan Quang. Nonparametric estimators for variable circular plot surveys. Biometrics, 49(3):837–852, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Evans:1993:CCJ

[124] Marc A. Evans and Douglas G. Bonett. A constrained Cook–Jacobson model of visibility bias. Biometrics, 49(3):853–859, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wellek:1993:LRT

[125] Stefan Wellek. A log-rank test for equivalence of two survivor functions. Biometrics, 49(3):877–881, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tomizawa:1993:DPS

[126] Sadao Tomizawa. Diagonals-parameter symmetry model for cumulative probabilities in square contingency tables with ordered categories. Biometrics, 49(3):883–887, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mendell:1993:WLR

[127] Nancy R. Mendell, Stephen J. Finch, and Henry C. Thode, Jr. Where is the likelihood ratio test powerful for detecting two component normal mixtures? Biometrics, 49(3):907–915, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kanji:1993:RFE

[128] G. K. Kanji. Review: Fundamentals of Exploratory Analysis of Variance., by D. C. Hoaglin, F. Mosteller, and J. W. Tukey. Biometrics, 49(3):958, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL http://www.jstor.org/stable/2532225.

Abt:1993:OAL

[129] Klaus Abt. Obituary: Arthur Linder 1904–1993. Biometrics, 49(3):961– 966, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lefkopoulou:1993:GTM

[130] Myrto Lefkopoulou and Louise Ryan. Global tests for multiple binary outcomes. Biometrics, 49(4):975–988, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dawson:1993:SPT

[131] Jeffrey D. Dawson and Stephen W. Lagakos. Size and power of twosample tests of repeated measures data. Biometrics, 49(4):1022–1032, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nandram:1993:BPI

[132] B. Nandram and J. Sedransk. Bayesian predictive inference for longitudinal sample surveys. Biometrics, 49(4):1045–1055, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hughes:1993:RDP

[133] Michael D. Hughes. Regression dilution in the proportional hazards model. Biometrics, 49(4):1056–1066, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wallenstein:1993:PMH

[134] Sylvan Wallenstein and Janet Wittes. The power of the Mantel–Haenszel test for grouped failure time data. Biometrics, 49(4):1077–1087, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Iwaisaki:1993:RGP

[135] Hiroaki Iwaisaki and James W. Wilton. Regression of genotypic on phenotypic value of a ratio-defined character. Biometrics, 49(4):1154–1163, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muttlak:1993:NLI

[136] Hassen A. Muttlak and S. M. Sadooghi-Alvandi. A note on the line intercept sampling method. Biometrics, 49(4):1209–1215, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Le:1993:TLT

[137] Chap T. Le. A test for linear trend in constant hazards and its application to a problem in occupational health. Biometrics, 49(4):1220–1224, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Becher:1993:BHT

[138] Heiko Becher. Bootstrap hypothesis testing procedures. Comment on: “Two guidelines for bootstrap hypothesis testing” [Biometrics 47 (1991), no. 2, 757–762; MR1132543 (92f:62061)] by P. G. Hall and S. R. Wilson with a reply by Hall and Wilson. Biometrics, 49(4):1268–1272, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [65].

Utts:1993:OFN

[139] Jessica Utts. Obituary: Florence Nightingale David (1909–1993). Biometrics, 49(4):1289–1291, 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Slud:1994:AFS

[140] Eric V. Slud. Analysis of factorial survival experiments. Biometrics, 50 (1):25–38, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muller:1994:HRE

[141] Hans-Georg Müller and Jane-Ling Wang. Hazard rate estimation under random censoring with varying kernels and bandwidths. Biometrics, 50 (1):61–76, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williamson:1994:MTD

[142] G. David Williamson and Michael Haber. Models for three-dimensional contingency tables with completely and partially cross-classified data. Biometrics, 50(1):194–203, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:1994:MPC

[143] Dean A. Follmann, Michael A. Proschan, and Nancy L. Geller. Monitoring pairwise comparisons in multi-armed clinical trials. Biometrics, 50(2): 325–336, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:1994:PBG

[144] Peter F. Thall and Richard Simon. Practical Bayesian guidelines for phase IIB clinical trials. Biometrics, 50(2):337–349, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Loisel:1994:DMG

[145] Patrice Loisel, Bruno Goffinet, Hervé Monod, and Gustavo Montes De Oca. Detecting a major gene in an F 2 population. Biometrics, 50(2): 512–516, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fitzmaurice:1994:AIL

[146] Garrett M. Fitzmaurice, Nan M. Laird, and Stuart R. Lipsitz. Analysing incomplete longitudinal binary responses: a likelihood-based approach. Biometrics, 50(3):601–612, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Commenges:1994:THB

[147] Daniel Commenges, Luc Letenneur, Hélène Jacqmin, Thierry Moreau, and Jean-François Dartigues. Test of homogeneity of binary data with explanatory variables. Biometrics, 50(3):613–620, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gray:1994:SBT

[148] Robert J. Gray. Spline-based tests in survival analysis. Biometrics, 50(3): 640–652, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kelly:1994:TMM

[149] Colleen Kelly. A test of the Markovian model of DNA evolution. Biometrics, 50(3):653–664, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Satten:1994:EIH

[150] Glen A. Satten and Ira M. Longini, Jr. Estimation of incidence of HIV infection using cross-sectional marker surveys. Biometrics, 50(3):675–688, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cressie:1994:SSO

[151] Noel Cressie and Jonathan Biele. A sample-size-optimal Bayesian procedure for sequential pharmaceutical trials. Biometrics, 50(3):700–711, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Link:1994:DEU

[152] William A. Link and Richard J. Barker. Density estimation using the trapping web design: a geometric analysis. Biometrics, 50(3):733–745, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shao:1994:SIS

[153] Jun Shao and Shein-Chung Chow. Statistical inference in stability analysis. Biometrics, 50(3):753–763, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Viana:1994:CML

[154] Marlos A. G. Viana. Combined maximum likelihood estimates for the equicorrelation coefficient. Biometrics, 50(3):813–820, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Silvapulle:1994:TAO

[155] Mervyn J. Silvapulle. On tests against one-sided hypotheses in some generalized linear models. Biometrics, 50(3):853–858, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Aksland:1994:GCA

[156] Magnar Aksland. A general cohort analysis method. Biometrics, 50(4): 917–932, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Conaway:1994:CNM

[157] Mark R. Conaway. Causal nonresponse models for repeated categorical measurements. Biometrics, 50(4):1102–1116, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hargrove:1994:PPP

[158] J. W. Hargrove and C. H. Borland. Pooled population parameter estimates from mark-recapture data. Biometrics, 50(4):1129–1141, 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thorne:1995:ERM

[159] Jeffrey L. Thorne and Gary A. Churchill. Estimation and reliability of molecular sequence alignments. Biometrics, 51(1):100–113, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hwang:1995:QEU

[160] W. D. Hwang and Anne Chao. Quantifying the effects of unequal catchabilities on Jolly–Seber estimators via sample coverage. Biometrics, 51(1): 128–141, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cox:1995:ACB

[161] Christopher Cox and Guangqin Ma. Asymptotic confidence bands for generalized nonlinear regression models. Biometrics, 51(1):142–150, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:1995:AGL

[162] Dean Follmann and Margaret Wu. An approximate generalized linear model with random effects for informative missing data. Biometrics, 51 (1):151–168, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barnhart:1995:MPM

[163] Huiman X. Barnhart and Allan R. Sampson. Multiple population models for multivariate random length data — with applications in clinical trials. Biometrics, 51(1):195–204, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:1995:SCI

[164] Shili Lin. A scheme for constructing an irreducible Markov chain for pedigree data. Biometrics, 51(1):318–322, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Andersen:1995:NTC

[165] Per Kragh Andersen and Birgitte B. Rønn. A nonparametric test for comparing two samples where all observations are either left- or rightcensored. Biometrics, 51(1):323–329, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zucker:1995:IAB

[166] David M. Zucker, Gary O. Zerbe, and Margaret C. Wu. Inference for the association between coefficients in a multivariate growth curve model. Biometrics, 51(2):413–424, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:1995:EMM

[167] Fraser B. Smith and Ronald W. Helms. EM mixed model analysis of data from informatively censored normal distributions. Biometrics, 51(2): 425–436, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rivest:1995:SMA

[168] Louis-Paul Rivest, François Potvin, Hélène Crépeau, and Gaétan Daigle. Statistical methods for aerial surveys using the double-count technique to correct visibility bias. Biometrics, 51(2):461–470, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

George:1995:FLP

[169] E. Olusegun George and Dale Bowman. A full likelihood procedure for analysing exchangeable binary data. Biometrics, 51(2):512–523, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

deBEdwardes:1995:CIE

[170] Michael D. deB. Edwardes. A confidence interval for Pr(X < Y)−Pr(X > Y) estimated from simple cluster samples. Biometrics, 51(2):571–578, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barry:1995:BMG

[171] Daniel Barry. A Bayesian model for growth curve analysis. Biometrics, 51 (2):639–655, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vounatsou:1995:BAR

[172] P. Vounatsou and A. F. M. Smith. Bayesian analysis of ring-recovery data via Markov chain Monte Carlo simulation. Biometrics, 51(2):687–708, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Drescher:1995:ACE

[173] Karsten Drescher and Wolfgang Boedeker. Assessment of the combined effects of substances: the relationship between concentration addition and independent action. Biometrics, 51(2):716–730, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

vandenBroek:1995:STZ

[174] Jan van den Broek. A score test for zero inflation in a Poisson distribution. Biometrics, 51(2):738–743, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Edwards:1995:XFM

[175] A. W. F. Edwards. XVIIIth Fisher Memorial Lecture delivered at the Natural History Museum, London, on Thursday, 20th October, 1994 — fiducial inference and the fundamental theorem of natural selection. Biometrics, 51(3):799–809, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lemdani:1995:TGL

[176] Mohamed Lemdani and Odile Pons. Tests for genetic linkage and homogeneity. Biometrics, 51(3):1033–1041, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maller:1995:TPI

[177] R. A. Maller and Sean Zhou. Testing for the presence of immune or cured individuals in censored survival data. Biometrics, 51(4):1197–1205, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sanchez:1995:STE

[178] Alex Sanchez, Jordi Ocaña, and Frederic Utzet. Sampling theory, estimation, and significance testing for Prevosti’s estimate of genetic distance. Biometrics, 51(4):1216–1235, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Singh:1995:PDC

[179] M. Singh and K. Hinkelmann. Partial diallel crosses in incomplete blocks. Biometrics, 51(4):1302–1314, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Karunamuni:1995:BEA

[180] Rohana J. Karunamuni and Terrance J. Quinn II. Bayesian estimation of animal abundance for line transect sampling. Biometrics, 51(4):1325– 1337, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bryant:1995:ITC

[181] John Bryant and Roger Day. Incorporating toxicity considerations into the design of two-stage phase II clinical trials. Biometrics, 51(4):1372–1383, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shih:1995:IAP

[182] Joanna H. Shih and Thomas A. Louis. Inferences on the association parameter in copula models for bivariate survival data. Biometrics, 51(4): 1384–1399, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Etchison:1995:PTS

[183] Tonya Etchison, Cavell Brownie, and Sastry G. Pantula. A portmanteau test for spatial ARMA models. Biometrics, 51(4):1536–1542, 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Christensen:1996:ETV

[184] Ronald Christensen. Exact tests for variance components. Biometrics, 52 (1):309–314, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheung:1996:SCI

[185] Siu Hung Cheung and Wai Sum Chan. Simultaneous confidence intervals for pairwise multiple comparisons in a two-way unbalanced design. Biometrics, 52(2):463–472, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

TenHave:1996:MEM

[186] Thomas R. Ten Have. A mixed effects model for multivariate ordinal response data including correlated discrete failure times with ordinal responses. Biometrics, 52(2):473–491, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hawkins:1996:ETP

[187] D. L. Hawkins, C. P. Han, and J. Eisenfeld. Estimating transition probabilities from aggregate samples augmented by haphazard recaptures. Biometrics, 52(2):625–638, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sammel:1996:LVM

[188] Mary Dupuis Sammel and Louise M. Ryan. Latent variable models with fixed effects. Biometrics, 52(2):650–663, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pradel:1996:UCM

[189] R. Pradel. Utilization of capture-mark-recapture for the study of recruitment and population growth rate. Biometrics, 52(2):703–709, 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cadima:1996:SSR

[190] Jorge F. C. L. Cadima and Ian T. Jolliffe. Size- and shape-related principal component analysis. Biometrics, 52(2):710–716, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qu:1996:REM

[191] Yinsheng Qu, Ming Tan, and Michael H. Kutner. Random effects models in latent class analysis for evaluating accuracy of diagnostic tests. Biometrics, 52(3):797–810, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fay:1996:RIT

[192] Michael P. Fay. Rank invariant tests for interval censored data under the grouped continuous model. Biometrics, 52(3):811–822, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Keen:1996:LES

[193] K. J. Keen. Limiting the effects of single-member families in the estimation of the intraclass correlation. Biometrics, 52(3):823–832, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mahdi:1996:CCS

[194] Smail Mahdi and Sabin Lessard. Convergence of covariance structures in additive Gaussian polygenic models. Biometrics, 52(3):833–845, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Troendle:1996:PSM

[195] James F. Troendle. A permutational step-up method of testing multiple outcomes. Biometrics, 52(3):846–859, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schwarz:1996:GMA

[196] Carl James Schwarz and A. Neil Arnason. A general methodology for the analysis of capture-recapture experiments in open populations. Biometrics, 52(3):860–873, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Quantin:1996:RSM

[197] Catherine Quantin, Thierry Moreau, Bernard Asselain, Jean Maccario, and Joseph Lellouch. A regression survival model for testing the proportional hazards hypothesis. Biometrics, 52(3):874–885, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cade:1996:PTL

[198] Brian S. Cade and Jon D. Richards. Permutation tests for least absolute deviation regression. Biometrics, 52(3):886–902, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Akritas:1996:NIF

[199] Michael G. Akritas and Michael P. LaValley. Nonparametric inference in factorial designs with censored data. Biometrics, 52(3):913–924, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ejigou:1996:PSS

[200] Ayenew Ejigou. Power and sample size for matched case-control studies. Biometrics, 52(3):925–933, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Walker:1996:EAN

[201] Stephen Walker. An EM algorithm for nonlinear random effects models. Biometrics, 52(3):934–944, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lauter:1996:ETA

[202] Jürgen Läuter. Exact t and F tests for analyzing studies with multiple endpoints. Biometrics, 52(3):964–970, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hirji:1996:NEA

[203] Karim F. Hirji. A note on exact analysis of several 2×2 tables. Biometrics, 52(3):1018–1025, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:1996:GSD

[204] D. Y. Lin, L. Shen, Z. Ying, and N. E. Breslow. Group sequential designs for monitoring survival probabilities. Biometrics, 52(3):1033–1041, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:1996:EMT

[205] Woollcott Smith and Andrew R. Solow. An exact McNemar test for paired binary Markov chains. Biometrics, 52(3):1063–1070, 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McKeown:1996:TAE

[206] Sean P. McKeown and William D. Johnson. Testing for autocorrelation and equality of covariance matrices. Biometrics, 52(3):1087–1095, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gatto:1996:SAM

[207] Riccardo Gatto. Saddlepoint approximations of marginal densities and confidence intervals in the logistic regression measurement error model. Biometrics, 52(3):1096–1102, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hepworth:1996:ECI

[208] Graham Hepworth. Exact confidence intervals for proportions estimated by group testing. Biometrics, 52(3):1134–1146, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Best:1996:NAD

[209] D. J. Best and J. C. W. Rayner. Nonparametric analysis for doubly ordered two-way contingency tables. Biometrics, 52(3):1153–1156, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bock:1996:HDM

[210] R. Darrell Bock and Robert D. Gibbons. High-dimensional multivariate probit analysis. Biometrics, 52(4):1183–1194, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:1996:MHT

[211] I-Ming Liu and Alan Agresti. Mantel–Haenszel-type inference for cumulative odds ratios with a stratified ordinal response. Biometrics, 52(4): 1223–1234, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fung:1996:DIO

[212] Wing K. Fung. Diagnosing influential observations in quadratic discriminant analysis. Biometrics, 52(4):1235–1241, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coakley:1996:VST

[213] Clint W. Coakley and Mark A. Heise. Versions of the sign test in the presence of ties. Biometrics, 52(4):1242–1251, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cierco:1996:CCI

[214] Christine Cierco and Brigitte Mangin. Construction of confidence intervals for quantitative trait loci location in F_{2} populations. Biometrics, 52(4): 1268–1282, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steele:1996:MEA

[215] Brian M. Steele. A modified EM algorithm for estimation in generalized mixed models. Biometrics, 52(4):1295–1310, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Akritas:1996:UNR

[216] Michael G. Akritas. On the use of nonparametric regression techniques for fitting parametric regression models. Biometrics, 52(4):1342–1362, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:1996:SSA

[217] Guohua Pan. Subset selection with additional order information. Biometrics, 52(4):1363–1374, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Conaway:1996:DPI

[218] Mark R. Conaway and Gina R. Petroni. Designs for Phase II trials allowing for a trade-off between response and toxicity. Biometrics, 52(4):1375–1386, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lewy:1996:GDD

[219] Peter Lewy. A generalized Dirichlet distribution accounting for singularities of the variables. Biometrics, 52(4):1394–1409, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shih:1996:TIB

[220] Joanna H. Shih and Thomas A. Louis. Tests of independence for bivariate survival data. Biometrics, 52(4):1440–1449, 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Langholz:1997:EAR

[221] B. Langholz and Ø. Borgan. Estimation of absolute risk from nested casecontrol data. Biometrics, 53(??):767–774, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [470].

Hung:1997:BVW

[222] H. M. James Hung, Robert T. O’Neill, Peter Bauer, and Karl Köhne. The behavior of the P -value when the alternative hypothesis is true. Biometrics, 53(1):11–22, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bar-Hen:1997:TSC

[223] Avner Bar-Hen and J.-J. Daudin. A test of a special case of typicality in linear discriminant analysis. Biometrics, 53(1):39–48, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chan:1997:MLE

[224] Jennifer S. K. Chan and Anthony Y. C. Kuk. Maximum likelihood estimation for probit-linear mixed models with correlated random effects. Biometrics, 53(1):86–97, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1997:RCF

[225] C. Y. Wang, Li Hsu, Z. D. Feng, and Ross L. Prentice. Regression calibration in failure time regression. Biometrics, 53(1):131–145, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eckert:1997:TAM

[226] R. Stephen Eckert, Raymond J. Carroll, and Naisyin Wang. Transformations to additivity in measurement error models. Biometrics, 53(1): 262–272, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Garcia-Soidan:1997:SRM

[227] Pilar H. Garcia-Soidan and Peter Hall. On sample reuse methods for spatial data. Biometrics, 53(1):273–281, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wulfsohn:1997:JMS

[228] Michael S. Wulfsohn and Anastasios A. Tsiatis. A joint model for survival and longitudinal data measured with error. Biometrics, 53(1):330–339, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Piepho:1997:DFT

[229] Hans-Peter Piepho. Distribution-free tests for one-way homoscedasticity in a two-way classification. Biometrics, 53(1):340–351, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Betensky:1997:ESA

[230] Rebecca A. Betensky. Early stopping to accept H_{0} based on conditional power: approximations and comparisons. Biometrics, 53(3):794– 806, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:1997:LMM

[231] Xihong Lin, Jonathan Raz, and Sioḃán D. Harlow. Linear mixed models with heterogeneous within-cluster variances. Biometrics, 53(3):910–923, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Waller:1997:LLM

[232] Lance A. Waller and Daniel Zelterman. Log-linear modeling with the negative multinomial distribution. Biometrics, 53(3):971–982, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Le:1997:ECE

[233] Chap T. Le. Evaluation of confounding effects in ROC studies. Biometrics, 53(3):998–1007, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lyles:1997:DEA

[234] Robert H. Lyles and Lawrence L. Kupper. A detailed evaluation of adjustment methods for multiplicative measurement error in linear regression with applications in occupational epidemiology. Biometrics, 53(3): 1008–1025, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gautam:1997:TLT

[235] Shiva Gautam. Test for linear trend in 2 × K ordered tables with openended categories. Biometrics, 53(3):1163–1169, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:1997:MHT

[236] Jie Zhang and Dennis D. Boos. Mantel–Haenszel test statistics for correlated binary data. Biometrics, 53(4):1185–1198, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Younes:1997:LBM

[237] Naji Younes and John Lachin. Link-based models for survival data with interval and continuous time censoring. Biometrics, 53(4):1199–1211, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Banneheka:1997:STS

[238] Sarath G. Banneheka, Richard D. Routledge, and Carl J. Schwarz. Stratified two-sample tag-recovery census of closed populations. Biometrics, 53 (4):1212–1224, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hougaard:1997:AOC

[239] Philip Hougaard, Mei-Ling Ting Lee, and G. A. Whitmore. Analysis of overdispersed count data by mixtures of Poisson variables and Poisson processes. Biometrics, 53(4):1225–1238, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sasieni:1997:GGD

[240] Peter D. Sasieni. From genotypes to genes: doubling the sample size. Biometrics, 53(4):1253–1261, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zeng:1997:CIB

[241] Qi Zeng and Marie Davidian. Calibration inference based on multiple runs of an immunoassay. Biometrics, 53(4):1304–1317, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burch:1997:ECI

[242] Brent D. Burch and Hari K. Iyer. Exact confidence intervals for a variance ratio (or heritability) in a mixed linear model. Biometrics, 53(4): 1318–1333, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Staniswalis:1997:SRA

[243] Joan G. Staniswalis, Peter F. Thall, and John Salch. Semiparametric regression analysis for recurrent event interval counts. Biometrics, 53(4): 1334–1353, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:1997:MDS

[244] Tiee-Jian Wu, John P. Burke, and Daniel B. Davison. A measure of DNA sequence dissimilarity based on Mahalanobis distance between frequencies of words. Biometrics, 53(4):1431–1439, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Andersen:1997:EVC

[245] Per Kragh Andersen, John P. Klein, Kim M. Knudsen, and René Tabanera y Palacios. Estimation of variance in Cox’s regression model with shared gamma frailties. Biometrics, 53(4):1475–1484, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:1997:SRS

[246] Wei Liu. Some results on step-up tests for comparing treatments with a control in unbalanced one-way layouts. Biometrics, 53(4):1508–1512, 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cullis:1998:SAM

[247] Brian Cullis, Bev Gogel, Arūnas Verbyla, and Robin Thompson. Spatial analysis of multi-environment early generation variety trials. Biometrics, 54(1):1–18, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Plante:1998:SCR

[248] Nathalie Plante, Louis-Paul Rivest, and Gilles Tremblay. Stratified capture-recapture estimation of the size of a closed population. Biometrics, 54(1):47–60, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:1998:PCI

[249] S. C. Cheng, Jason P. Fine, and L. J. Wei. Prediction of cumulative incidence function under the proportional hazards model. Biometrics, 54 (1):219–228, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stallard:1998:SSD

[250] Nigel Stallard. Sample size determination for Phase II clinical trials based on Bayesian decision theory. Biometrics, 54(1):279–294, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Olkin:1998:CMA

[251] Ingram Olkin and Allan Sampson. Comparison of meta-analysis versus analysis of variance of individual patient data. Biometrics, 54(1):317–322, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:1998:CCA

[252] Xiao-Hua Zhou. Comparing correlated areas under the ROC curves of two diagnostic tests in the presence of verification bias. Biometrics, 54(2): 453–470, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

StLaurent:1998:EAG

[253] Roy T. St. Laurent. Evaluating agreement with a gold standard in method comparison studies. Biometrics, 54(2):537–545, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mack:1998:KML

[254] Y. P. Mack and Pham X. Quang. Kernel methods in line and point transect sampling. Biometrics, 54(2):606–619, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Walker:1998:NAS

[255] Stephen Walker. A nonparametric approach to a survival study with surrogate endpoints. Biometrics, 54(2):662–672, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hommel:1998:IMT

[256] Gerhard Hommel and Frank Krummenauer. Improvements and modifications of Tarone’s multiple test procedure for discrete data. Biometrics, 54 (2):673–681, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:1998:MEL

[257] Song Xi Chen. Measurement errors in line transect surveys. Biometrics, 54 (3):899–908, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

FitzGerald:1998:IRL

[258] Patrick E. B. FitzGerald and Matthew W. Knuiman. Interpretation of regressive logistic regression coefficients in analyses of familial data. Biometrics, 54(3):909–920, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gray:1998:ETE

[259] Sarah M. Gray and Ron Brookmeyer. Estimating a treatment effect from multidimensional longitudinal data. Biometrics, 54(3):976–988, 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Martinussen:1998:ADU

[260] Torben Martinussen and Thorkild I. A. Sørensen. Age-dependent Ushaped risk functions and Aalen’s additive risk model. Biometrics, 54(3): 989–1001, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lipsitz:1998:EEI

[261] Stuart R. Lipsitz and Joseph G. Ibrahim. Estimating equations with incomplete categorical covariates in the Cox model. Biometrics, 54(3): 1002–1013, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:1998:SSE

[262] Song Yang. Some scale estimators and lack-of-fit tests for the censored two-sample accelerated life model. Biometrics, 54(3):1040–1052, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shih:1998:MMD

[263] Joanna H. Shih. Modeling multivariate discrete failure time data. Biometrics, 54(3):1115–1128, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:1998:ODS

[264] Qing Liu. An order-directed score test for trend in ordered 2 × K tables. Biometrics, 54(3):1147–1154, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leon:1998:AMA

[265] Larry F. Léon and Chih-Ling Tsai. Assessment of model adequacy for Markov regression time series models. Biometrics, 54(3):1165–1175, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Teng:1998:MLA

[266] Jun Teng and David Siegmund. Multipoint linkage analysis using affected relative pairs and partially informative markers. Biometrics, 54(4): 1247–1279, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). With discussion and a rejoinder by the authors.

Hu:1998:EPC

[267] Ping Hu, Anastasios A. Tsiatis, and Marie Davidian. Estimating the parameters in the Cox model when covariate variables are measured with error. Biometrics, 54(4):1407–1419, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:1998:NRE

[268] Haesook T. Kim and Young K. Truong. Nonparametric regression estimates with censored data: local linear smoothers and their applications.

Biometrics, 54(4):1434–1444, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Faraggi:1998:BVS

[269] David Faraggi and Richard Simon. Bayesian variable selection method for censored survival data. Biometrics, 54(4):1475–1485, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Verweij:1998:GFT

[270] Pierre J. M. Verweij, Hans C. van Houwelingen, and Theo Stijnen. A goodness-of-fit test for Cox’s proportional hazards model based on martingale residuals. Biometrics, 54(4):1517–1526, 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mau:1999:BPI

[271] Bob Mau, Michael A. Newton, and Bret Larget. Bayesian phylogenetic inference via Markov chain Monte Carlo methods. Biometrics, 55(1):1– 12, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hougaard:1999:FSD

[272] Philip Hougaard. Fundamentals of survival data. Biometrics, 55(1):13– 22, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Keiding:1999:GVL

[273] Niels Keiding, Marusca Filiberti, Sille Esbjerg, James M. Robins, and Niels Jacobsen. The graft versus leukemia effect after bone marrow transplantation: a case study using structural nested failure time models. Biometrics, 55(1):23–28, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Skaug:1999:HML

[274] Hans J. Skaug and Tore Schweder. Hazard models for line transect surveys with independent observers. Biometrics, 55(1):29–36, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robin:1999:MT

[275] Stéphane Robin. A model for thermograms. Biometrics, 55(1):37–43, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carroll:1999:FPM

[276] Raymond J. Carroll, Kathryn Roeder, and Larry Wasserman. Flexible parametric measurement error models. Biometrics, 55(1):44–54, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Darling:1999:GMR

[277] R. W. R. Darling and Tim Holt. Genetic models with reduced penetrance related to the Y chromosome. Biometrics, 55(1):55–64, March 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liang:1999:HTU

[278] Kung-Yee Liang and Paul J. Rathouz. Hypothesis testing under mixture models: Application to genetic linkage analysis. Biometrics, 55(1):65– 74, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:1999:USM

[279] Margaret C. Wu and Dean A. Follmann. Use of summary measures to adjust for informative missingness in repeated measures data with random effects. Biometrics, 55(1):75–84, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

TenHave:1999:MEM

[280] Thomas R. Ten Have and Alfredo Morabia. Mixed effects models with bivariate and univariate association parameters for longitudinal bivariate binary response data. Biometrics, 55(1):85–93, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Golm:1999:SMM

[281] Gregory T. Golm, M. Elizabeth Halloran, and Ira M. Longini, Jr. Semiparametric methods for multiple exposure mismeasurement and a bivariate outcome in HIV vaccine trials. Biometrics, 55(1):94–101, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tao:1999:EMS

[282] Huageng Tao, Mari Palta, Brian S. Yandell, and Michael A. Newton. An estimation method for the semiparametric mixed effects model. Biometrics, 55(1):102–110, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1999:BCC

[283] Jinping Wang and Sabyasachi Basu. Bias-corrected confidence intervals for the concentration parameter in a dilution assay. Biometrics, 55(1):111– 116, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Aitkin:1999:GML

[284] Murray Aitkin. A general maximum likelihood analysis of variance components in generalized linear models. Biometrics, 55(1):117–128, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vangel:1999:MLA

[285] Mark G. Vangel and Andrew L. Rukhin. Maximum likelihood analysis for heteroscedastic one-way random effects ANOVA in interlaboratory studies. Biometrics, 55(1):129–136, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:1999:EIC

[286] Martin S. Ridout, Clarice G. B. Demétrio, and David Firth. Estimating intraclass correlation for binary data. Biometrics, 55(1):137–148, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [469].

Lindsey:1999:RSO

[287] J. K. Lindsey. Response surfaces for overdispersion in the study of the conditions for fish eggs hatching. Biometrics, 55(1):149–155, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gignoux:1999:CPD

[288] Jacques Gignoux, Camille Duby, and Sébastian Barot. Comparing the performances of Diggle’s tests of spatial randomness for small samples with and without edge-effect correction: Application to ecological data. Biometrics, 55(1):156–164, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greenhalgh:1999:PRM

[289] David Greenhalgh, Murray T. Doyle, and Janet Mortimer. A partial ranking method for identifying repeated inclusion of individuals in anonymized HIV infection reports. Biometrics, 55(1):165–173, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Do:1999:DAE

[290] Kim-Anh Do and Katherine Kirk. Discriminant analysis of event-related potential curves using smoothed principal components. Biometrics, 55 (1):174–181, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:1999:REM

[291] Jianwen Cai, Pranab K. Sen, and Haibo Zhou. A random effects model for multivariate failure time data from multicenter clinical trials. Biometrics, 55(1):182–189, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:1999:SIS

[292] Yu Shen and Lloyd Fisher. Statistical inference for self-designing clinical trials with a one-sided hypothesis. Biometrics, 55(1):190–197, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huzurbazar:1999:SHF

[293] S. Huzurbazar and Aparna V. Huzurbazar. Survival and hazard functions for progressive diseases using saddlepoint approximations. Biometrics, 55 (1):198–203, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

LeBlanc:1999:ARS

[294] Michael LeBlanc and John Crowley. Adaptive regression splines in the Cox model. Biometrics, 55(1):204–213, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lipsitz:1999:LMI

[295] Stuart R. Lipsitz, Joseph G. Ibrahim, and Garrett M. Fitzmaurice. Likelihood methods for incomplete longitudinal binary responses with incomplete categorical covariates. Biometrics, 55(1):214–223, March 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Routledge:1999:EST

[296] R. D. Routledge, G. E. J. Smith, L. Sun, N. Dawe, E. Nygren, and J. S. Sedinger. Estimating the size of a transient population. Biometrics, 55 (1):224–230, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hung:1999:RGT

[297] M. Hung and William H. Swallow. Robustness of group testing in the estimation of proportions. Biometrics, 55(1):231–237, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Betensky:1999:LEE

[298] Rebecca A. Betensky, Jane C. Lindsey, Louise M. Ryan, and M. P. Wand. Local EM estimation of the hazard function for interval-censored data. Biometrics, 55(1):238–245, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kolassa:1999:AMC

[299] John E. Kolassa and Martin A. Tanner. Approximate Monte Carlo conditional inference in exponential families. Biometrics, 55(1):246–251, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:1999:PRM

[300] D. Y. Lin, P. Arbogast, D. S. Siscovick, and R. N. Lemaitre. Poisson regression with missing durations of exposure. Biometrics, 55(1):252–257, 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ryan:1999:PRM

[301] Louise M. Ryan, D. Y. Lin, P. Arbogast, D. S. Siscovick, and R. N. Lemaitre. Poisson regression with missing durations of exposure. Biometrics, 55(1):252–257, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tan:1999:RLV

[302] Ming Tan, Yinsheng Qu, and J. Sunil Rao. Robustness of the latent variable model for correlated binary data. Biometrics, 55(1):258–263, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Freidlin:1999:UVS

[303] Boris Freidlin and Joseph L. Gastwirth. Unconditional versions of several tests commonly used in the analysis of contingency tables. Biometrics, 55(1):264–267, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liao:1999:HBM

[304] J. G. Liao. A hierarchical Bayesian model for combining multiple 2 × 2 tables using conditional likelihoods. Biometrics, 55(1):268–272, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shukla:1999:NET

[305] G. K. Shukla and G. S. V. Subrahmanyam. A note on an exact test and confidence interval for competition and overlap effects. Biometrics, 55 (1):273–276, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:1999:GST

[306] Zhengqing Li. A group sequential test for survival trials: an alternative to rank-based procedures. Biometrics, 55(1):277–283, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:1999:UMF

[307] Richard J. Cook and Vernon T. Farewell. The utility of mixed-form likelihoods. Biometrics, 55(1):284–288, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nam:1999:PSS

[308] Jun-Mo Nam. Power and sample size for testing homogeneity of relative risks in prospective studies. Biometrics, 55(1):289–293, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coull:1999:UML

[309] Brent A. Coull and Alan Agresti. The use of mixed logit models to reflect heterogeneity in capture–recapture studies. Biometrics, 55(1):294– 301, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Begg:1999:ATU

[310] Melissa Dowd Begg. Analyzing k(2 × 2) tables under cluster sampling. Biometrics, 55(1):302–307, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xie:1999:RMO

[311] Minge Xie and Douglas G. Simpson. Regression modeling of ordinal data with nonzero baselines. Biometrics, 55(1):308–316, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Betensky:1999:MSS

[312] Rebecca A. Betensky and Daniel Rabinowitz. Maximally selected x^{2} statistics for k×2 tables. Biometrics, 55(1):317–320, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bakke:1999:EOP

[313] Øyvind Bakke and Svein-Håkon Lorentsen. Estimation of offspring production from a limited number of stage-structured censuses. Biometrics, 55(1):321–325, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kemp:1999:BRa

[314] C. D. Kemp and A. W. Kemp. Book reviews. Biometrics, 55(1):326– 332, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wilson:1999:EB

[315] Susan R. Wilson. Evolution and biometry. Biometrics, 55(2):333–337, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morrissey:1999:MME

[316] Mary J. Morrissey and Donna Spiegelman. Matrix methods for estimating odds ratios with misclassified exposure data: Extensions and comparisons. Biometrics, 55(2):338–344, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cui:1999:NED

[317] Jisheng Cui. Nonparametric estimation of a delay distribution based on left-censored and right-truncated data. Biometrics, 55(2):345–349, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wei:1999:ODD

[318] Run-Peng Wei and Francis C. Yeh. Optimal diversity-dependent contributions of genotypes to mixtures. Biometrics, 55(2):350–354, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rongling:1999:MEM

[319] Wu Rongling and Bailian Li. A multiplicative–epistatic model for analyzing interspecific differences in outcrossing species. Biometrics, 55(2): 355–365, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stanley:1999:GFT

[320] Thomas R. Stanley and Kenneth P. Burnham. A goodness-of-fit test for capture–recapture model M_{t} under closure. Biometrics, 55(2):366– 375, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:1999:SDG

[321] Jae Kyun Lee, Martin Lascoux, Michael A. Newton, and Erik V. Nrdheim. A study of deleterious gene structure in plants using Markov chain Monte Carlo. Biometrics, 55(2):376–386, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huggins:1999:ESO

[322] Richard M. Huggins and Paul S. F. Yip. Estimation of the size of an open population from capture–recapture data using weighted martingale methods. Biometrics, 55(2):387–395, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sitter:1999:TSD

[323] R. R. Sitter and C. F. J. Wu. Two-stage design of quanta1 response studies. Biometrics, 55(2):396–402, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:1999:RPR

[324] Dean A. Follmann, Sally A. Hunsberger, and Paul S. Albert. Repeated probit regression when covariates are measured with error. Biometrics, 55(2):403–409, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:1999:PHD

[325] Hulin Wu and A. Adam Ding. Population HIV-1 dynamics in vivo: Applicable models and inferential tools for virological data from AIDS clinical trials. Biometrics, 55(2):410–418, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

George:1999:BAO

[326] A. W. George, K. L. Mengersen, and G. P. Davis. A Bayesian approach to ordering gene markers. Biometrics, 55(2):419–429, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barber:1999:STC

[327] Stuart Barber and Christopher Jennison. Symmetric tests and confidence intervals for survival probabilities and quantiles of censored survival data. Biometrics, 55(2):430–436, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zocchi:1999:OED

[328] Silvio S. Zocchi and Anthony C. Atkinson. Optimum experimental designs for multinomial logistic models. Biometrics, 55(2):437–444, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goggins:1999:ACP

[329] William B. Goggins, Dianne M. Finkelstein, and Alan M. Zaslavsky. Applying the Cox proportional hazards model when the change time of a binary time-varying covariate is interval censored. Biometrics, 55(2):445– 451, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:1999:AIG

[330] Heping Zhang. Analysis of infant growth curves using multivariate adaptive splines. Biometrics, 55(2):452–459, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Quan:1999:ART

[331] Hui Quan and Thomas Capizzi. Adjusted regression trend test for a multicenter clinical trial. Biometrics, 55(2):460–462, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muthen:1999:FMM

[332] Bengt Muthén and Kerby Shedden. Finite mixture modeling with mixture outcomes using the EM algorithm. Biometrics, 55(2):463–469, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berlin:1999:ECS

[333] Jesse A. Berlin, Stephen E. Kimmel, Thomas R. Ten Have, and Mary D. Sammel. An empirical comparison of several clustered data approaches under confounding due to cluster effects in the analysis of complications of coronary angioplasty. Biometrics, 55(2):470–476, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Walker:1999:BSA

[334] Stephen Walker and Bani K. Mallick. A Bayesian semiparametric accelerated failure time model. Biometrics, 55(2):477–483, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Simon:1999:BDA

[335] Richard Simon. Bayesian design and analysis of active control clinical trials. Biometrics, 55(2):484–487, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Toledano:1999:GEE

[336] Alicia Y. Toledano and Constantine Gatsonis. Generalized estimating equations for ordinal categorical data: Arbitrary patterns of missing responses and missingness in a key covariate. Biometrics, 55(2):488–496, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Klein:1999:MRE

[337] John P. Klein, Corey Pelz, and Mei jie Zhang. Modeling random effects for censored data by a multivariate normal regression model. Biometrics, 55(2):497–506, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Satten:1999:FSM

[338] Glen A. Satten and Maya R. Sternberg. Fitting semi-Markov models to interval-censored data with unknown initiation times. Biometrics, 55(2): 507–513, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sternberg:1999:DTN

[339] Maya R. Sternberg and Glen A. Satten. Discrete-time nonparametric estimation for semi-Markov models of chain-of-events data subject to interval censoring and truncation. Biometrics, 55(2):514–522, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nardi:1999:NRC

[340] Alessandra Nardi and Michael Schemper. New residuals for Cox regression and their application to outlier screening. Biometrics, 55(2):523–529, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

vanderLaan:1999:LEE

[341] Mark J. van der Laan and Alan Hubbard. Locally efficient estimation of the quality-adjusted lifetime distribution with right-censored data and covariates. Biometrics, 55(2):530–536, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:1999:SMS

[342] D. B. Dunson, C. R. Weinberg, S. D. Perreault, and R. E. Chapin. Summarizing the motion of self-propelled cells: Applications to sperm motility. Biometrics, 55(2):537–543, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kulldorff:1999:KMO

[343] Martin Kulldorff and Ulf Hjalmars. The Knox method and other tests for space–time interaction. Biometrics, 55(2):544–552, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Natarajan:1999:MHN

[344] Ranjini Natarajan and Charles E. McCulloch. Modeling heterogeneity in nest survival data. Biometrics, 55(2):553–559, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:1999:RSH

[345] Paul R. Rosenbaum. Reduced sensitivity to hidden bias at upper quantiles in observational studies with dilated treatment effects. Biometrics, 55(2): 560–564, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Deltour:1999:SAM

[346] Isabelle Deltour, Sylvia Richardson, and Jean-Yves Le Hesran. Stochastic algorithms for Markov models estimation with intermittent missing data. Biometrics, 55(2):565–573, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Preisser:1999:RRC

[347] John S. Preisser and Bahjat F. Qaqish. Robust regression for clustered data with application to binary responses. Biometrics, 55(2):574–579, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Parzen:1999:GGF

[348] Michael Parzen and Stuart R. Lipsitz. A global goodness-of-fit statistic for Cox regression models. Biometrics, 55(2):580–584, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinha:1999:BAM

[349] Debajyoti Sinha, Ming-Hui Chen, and Sujit K. Ghosh. Bayesian analysis and model selection for interval-censored survival data. Biometrics, 55(2):

585–590, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ibrahim:1999:MCE

[350] Joseph G. Ibrahim, Ming-Hui Chen, and Stuart R. Lipsitz. Monte Carlo EM for missing covariates in parametric regression models. Biometrics, 55(2):591–596, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agresti:1999:LCI

[351] Alan Agresti. On logit confidence intervals for the odds ratio with small samples. Biometrics, 55(2):597–602, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:1999:BME

[352] Dean A. Follmann and Paul S. Albert. Bayesian monitoring of event rates with censored data. Biometrics, 55(2):603–607, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [471].

Brookmeyer:1999:AMP

[353] Ron Brookmeyer. Analysis of multistage pooling studies of biological specimens for estimating disease incidence and prevalence. Biometrics, 55(2):608–612, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:1999:SVC

[354] Xihong Lin and Raymond J. Carroll. SIMEX variance component tests in generalized linear mixed measurement error models. Biometrics, 55(2): 613–619, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McDonald:1999:ETG

[355] John W. McDonald, Peter W. F. Smith, and Jonathan J. Forster. Exact tests of goodness of fit of log–linear models for rates. Biometrics, 55(2): 620–624, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hughes:1999:MEM

[356] James P. Hughes. Mixed effects models with censored data with application to HIV RNA levels. Biometrics, 55(2):625–629, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greenwood:1999:SIM

[357] Celia M. T. Greenwood, Julian P. Midgley, Andrew G. Matthew, and Alexander G. Logan. Statistical issues in a metaregression analysis of randomized trials: Impact on the dietary sodium intake and blood pressure relationship. Biometrics, 55(2):630–636, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qiou:1999:MSA

[358] Zuqiang Qiou, Nalini Ravishanker, and Dipak K. Dey. Multivariate survival analysis with positive stable frailties. Biometrics, 55(2):637–644, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xiao-Hua:1999:CSI

[359] Zhou Xiao-Hua and Wanzhu Tu. Comparison of several independent population means when their samples contain log–normal and possibly zero observations. Biometrics, 55(2):645–651, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Graubard:1999:PMS

[360] Barry I. Graubard and Edward L. Korn. Predictive margins with survey data. Biometrics, 55(2):652–659, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:1999:MCT

[361] Martin S. Ridout. Memory in coal tits: an alternative model. Biometrics, 55(2):660–662, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rayner:1999:MTS

[362] J. C. W. Rayner and D. J. Best. Modelling ties in the sign test. Biometrics, 55(2):663–665, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kemp:1999:BRb

[363] C. D. Kemp and A. W. Kemp. Book reviews. Biometrics, 55(2):666– 670, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stockmarr:1999:LRE

[364] Anders Stockmarr. Likelihood ratios for evaluating DNA evidence when the suspect is found through a database search. Biometrics, 55(3):671–677, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schweder:1999:SLM

[365] Tore Schweder, Hans J. Skaug, Mette Langaas, and Xeni K. Dimakos. Simulated likelihood methods for complex double-platform line transect surveys. Biometrics, 55(3):678–687, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heagerty:1999:MSL

[366] Patrick J. Heagerty. Marginally specified logistic-normal models for longitudinal binary data. Biometrics, 55(3):688–698, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Durban:1999:ASE

[367] Maria Durban, Christine A. Hackett, and I. D. Currie. Approximate standard errors in semiparametric models. Biometrics, 55(3):699–703, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Opsomer:1999:KNV

[368] J. D. Opsomer, D. Ruppert, M. P. Wand, U. Holst, and O. Hössjer. Kriging with nonparametric variance function estimation. Biometrics, 55(3):704– 710, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1999:ECB

[369] C. Y. Wang, Garnet L. Anderson, and Ross L. Prentice. Estimation of the correlation between nutrient intake measures under restricted sampling. Biometrics, 55(3):711–717, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Weinberg:1999:UPE

[370] Clarice R. Weinberg and David M. Umbach. Using pooled exposure assessment to improve efficiency in case-control studies. Biometrics, 55(3): 718–726, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [472].

Chavez-Demoulin:1999:BIS

[371] V. Chavez-Demoulin. Bayesian inference for small-sample capture– recapture data. Biometrics, 55(3):727–731, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:1999:VIR

[372] Dean A. Follmann and Michael A. Proschan. Valid inference in random effects meta-analysis. Biometrics, 55(3):732–737, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heikkinen:1999:MPF

[373] Juha Heikkinen and Elja Arjas. Modeling a Poisson forest in variable elevations: a nonparametric Bayesian approach. Biometrics, 55(3):738– 745, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:1999:TCB

[374] Peter F. Thall and Su-Chun Chengt. Treatment comparisons based on two-dimensional safety and efficacy alternatives in oncology trials. Biometrics, 55(3):746–753, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:1999:EIO

[375] Song Xi Chen. Estimation in independent observer line transect surveys for clustered populations. Biometrics, 55(3):754–759, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Regan:1999:LMC

[376] Meredith M. Regan and Paul J. Catalano. Likelihood models for clustered binary and continuous out comes: Application to developmental toxicology. Biometrics, 55(3):760–768, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gerard:1999:LBS

[377] Patrick D. Gerard and William R. Schucany. Local bandwidth selection for kernel estimation of population densities with line transect sampling. Biometrics, 55(3):769–773, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pennello:1999:UME

[378] Gene A. Pennello, Susan S. Devesa, and Mitchell H. Gail. Using a mixed effects model to estimate geographic variation in cancer rates. Biometrics, 55(3):774–781, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:1999:SPT

[379] Dean A. Follmann and Michael A. Proschan. A simple permutation-type method for testing circular uniformity with correlated angular measurements. Biometrics, 55(3):782–791, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Datta:1999:EEV

[380] Susmita Datta, M. Elizabeth Halloran, and Ira M. Longini, Jr. Efficiency of estimating vaccine efficacy for susceptibility and infectiousness: Randomization by individual versus household. Biometrics, 55(3):792–798, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:1999:CDB

[381] Song Xi Chen and Jodie L. Woolcock. A condition for designing bus-route type access site surveys to estimate recreational fishing effort. Biometrics, 55(3):799–804, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Clegg:1999:MMB

[382] Limin X. Clegg, Jianwen Cai, and Pranab K. Sen. A marginal mixed baseline hazards model for multivariate failure time data. Biometrics, 55 (3):805–812, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ross:1999:MCD

[383] Eric A. Ross and Dirk Moore. Modeling clustered, discrete, or grouped time survival data with covariates. Biometrics, 55(3):813–819, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dryden:1999:HRR

[384] Ian L. Dryden and Gary Walker. Highly resistant regression and object matching. Biometrics, 55(3):820–825, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cho:1999:FLA

[385] Meehyung Cho and Nathaniel Schenker. Fitting the log-F accelerated failure time model with incomplete covariate data. Biometrics, 55(3): 826–833, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rabinowitz:1999:TAO

[386] Daniel Rabinowitz and Qiong Yang. Testing for age-at-onset anticipation with affected parent–child pairs. Biometrics, 55(3):834–838, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reboussin:1999:EEL

[387] Beth A. Reboussin, Kung-Yee Liang, and David M. Reboussin. Estimating equations for a latent transit ion model with multiple discrete indicators. Biometrics, 55(3):839–845, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ofversten:1999:SQC

[388] Jukka Öfversten. Statistical quality control of variety trial data. Biometrics, 55(3):846–852, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cui:1999:MSS

[389] Lu Cui, H. M. James Hung, and Sue-Jane Wang. Modification of sample size in group sequential clinical trials. Biometrics, 55(3):853–857, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muller:1999:BHA

[390] Peter Muller, Giovanni Parmigiani, Joellen Schildkraut, and Luca Tardella. A Bayesian hierarchical approach for combining case-control and prospective studies. Biometrics, 55(3):858–866, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:1999:DS

[391] Wesley O. Johnson and Larry M. Pearson. Dual screening. Biometrics, 55 (3):867–873, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kittelson:1999:UFG

[392] John M. Kittelson and Scott S. Emerson. A unifying family of group sequential test designs. Biometrics, 55(3):874–882, September 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ryan:1999:ERT

[393] Louise M. Ryan, Boris Freidlin, Marvin J. Podgor, and Joseph L. Gastwirth. Efficiency robust tests for survival or ordered categorical data.

Biometrics, 55(3):883–886, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joly:1999:PLA

[394] Pierre Joly and Daniel Commenges. A penalized likelihood approach for a progressive three-state model with censored and truncated data: Application to AIDS. Biometrics, 55(3):887–890, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Real:1999:SEH

[395] Daniel Real and Ian L. Gordon. Standard errors of heritabilities for forage breeding nurseries. Biometrics, 55(3):891–895, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:1999:ADE

[396] Yanqing Sun, Shein-Chung Chow, Gang Li, and Keh-Wei Chen. Assessing distributions of estimated drug shelf lives in stability analysis. Biometrics, 55(3):896–899, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1999:EEP

[397] You-Gan Wang. Estimating equations for parameters in stochastic growth models from tag-recapture data. Biometrics, 55(3):900–903, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yip:1999:EPS

[398] Paul S. F. Yip, Yong Zhou, D. Y. Lin, and Xiang-Zhong Fang. Estimation of population size based on additive hazards models for continuous-time recapture experiments. Biometrics, 55(3):904–908, September 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:1999:RAD

[399] Jianguo Sun, Qiming Liao, and Marcello Pagano. Regression analysis of doubly censored failure time data with applications to AIDS studies. Biometrics, 55(3):909–914, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:1999:MMT

[400] Richard J. Cook. A mixed model for two-state Markov processes under panel observation. Biometrics, 55(3):915–920, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Weerahandi:1999:EIG

[401] Samaradasa Weerahandi and Vance W. Berger. Exact inference for growth curves with intraclass correlation structure. Biometrics, 55(3):921–924, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dette:1999:ODW

[402] Holger Dette and Weng Kee Wong. Optimal designs when the variance is a function of the mean. Biometrics, 55(3):925–929, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qiu:1999:MDS

[403] Peihua Qiu, Rick Chappell, William Obermeyer, and Ruth Benca. Modeling daily and subdaily cycles in rat sleep data. Biometrics, 55(3):930–935, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agresti:1999:MCV

[404] Alan Agresti and I-Ming Liu. Modeling a categorical variable allowing arbitrarily many category choices. Biometrics, 55(3):936–943, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pepe:1999:PUP

[405] Margaret Sullivan Pepe, Patrick Heagerty, and Robert Whitaker. Prediction using partly conditional time-varying coefficients regression models. Biometrics, 55(3):944–950, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fahrmeir:1999:AIL

[406] Ludwig Fahrmeir, Christian Gieger, and Christian Heumann. An application of isotonic longitudinal marginal regression to monitoring the healing process. Biometrics, 55(3):951–956, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mikulich:1999:RCC

[407] Susan K. Mikulich, Gary O. Zerbe, Richard H. Jones, and Thomas J. Crowley. Relating the classical covariance adjustment techniques of multivariate growth curve models to modern univariate mixed effects models. Biometrics, 55(3):957–964, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:1999:MTO

[408] D. B. Dunson and J. K. Haseman. Modeling tumor onset and multiplicity using transition models with latent variables. Biometrics, 55(3):965–970, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stallard:1999:DTD

[409] Nigel Stallard, Peter F. Thall, and John Whitehead. Decision theoretic designs for Phase II clinical trials with multiple outcomes. Biometrics, 55 (3):971–977, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Michiels:1999:SMP

[410] Bart Michiels, Geert Molenberghs, and Stuart R. Lipsitz. Selection models and pattern-mixture models for incomplete data with covariates. Biometrics, 55(3):978–983, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1999:EEN

[411] You-Gan Wang. Estimating equations with nonignorably missing response data. Biometrics, 55(3):984–989, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greenland:1999:LEB

[412] Sander Greenland, D. Y. Lin, R. A. Kronmal, and B. M. Psaty. Letter to the Editor of Biometrics. Biometrics, 55(3):990–991, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kemp:1999:BRc

[413] C. D. Kemp and A. W. Kemp. Book reviews. Biometrics, 55(3):992–996, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Devlin:1999:GCA

[414] B. Devlin and Kathryn Roeder. Genomic control for association studies. Biometrics, 55(4):997–1004, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Royston:1999:NAM

[415] Patrick Royston and Alberto Ferreira. A new approach to modeling daily probabilities of conception. Biometrics, 55(4):1005–1013, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schwarz:1999:SPE

[416] Carl J. Schwarz, Myra Andrews, and Michael R. Link. The stratified Petersen estimator with a known number of unread tags. Biometrics, 55(4):1014–1021, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

TenHave:1999:EBE

[417] Thomas R. Ten Have and A. Russell Localio. Empirical Bayes estimation of random effects parameters in mixed effects logistic regression models. Biometrics, 55(4):1022–1029, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Minkin:1999:LBE

[418] Salomon Minkin and Kiran Kundhal. Likelihood-based experimental design for estimation of Ed 50. Biometrics, 55(4):1030–1037, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:1999:EBE

[419] D. G. Chen, E. M. Carter, J. J. Hubert, and P. T. Kim. Empirical Bayes estimation for combinations of multivariate bioassays. Biometrics, 55(4): 1038–1043, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Halpern:1999:MSO

[420] Aaron L. Halpern. Minimally selected P and other tests for a single abrupt changepoint in a binary sequence. Biometrics, 55(4):1044–1050, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hogmander:1999:MSP

[421] Harri Högmander and Aila SäUrkkä. Multitype spatial point patterns with hierarchical interactions. Biometrics, 55(4):1051–1058, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reiczigel:1999:AED

[422] Jenö Reiczigel. Analysis of experimental data with repeated measurements. Biometrics, 55(4):1059–1063, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McGuire:1999:IEB

[423] Gráinne McGuire, Michael J. Prentice, and Frank Wright. Improved error bounds for genetic distances from DNA sequences. Biometrics, 55(4):

1064–1070, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mallick:1999:BSA

[424] Bani K. Mallick, David G. T. Denison, and Adrian F. M. Smith. Bayesian survival analysis using a Mars model. Biometrics, 55(4):1071–1077, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mahe:1999:ERP

[425] Cédric Mahé and Sylvie Chevret. Estimating regression parameters and degree of dependence for multivariate failure time data. Biometrics, 55(4): 1078–1084, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Murray:1999:SMC

[426] Susan Murray and Anastasios A. Tsiatis. Sequential methods for comparing years of life saved in the two-sample censored data problem. Biometrics, 55(4):1085–1092, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:1999:CBC

[427] Yu Shen and S. C. Cheng. Confidence bands for cumulative incidence curves under the additive risk model. Biometrics, 55(4):1093–1100, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:1999:EED

[428] Hongwei Zhao and Anastasios A. Tsiatis. Efficient estimation of the distribution of quality-adjusted survival time. Biometrics, 55(4):1101–1107, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:1999:TSP

[429] Yijian Huang. The two-sample problem with induced dependent censorship. Biometrics, 55(4):1108–1113, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chang:1999:NAD

[430] Chung-Chou H. Chang and Lisa A. Weissfeld. Normal approximation diagnostics for the Cox model. Biometrics, 55(4):1114–1119, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Piepho:1999:FRM

[431] Hans-Peter Piepho. Fitting a regression model for genotype-by-environment data on heading dates in grasses by methods for nonlinear mixed models. Biometrics, 55(4):1120–1128, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Langholz:1999:ABR

[432] Bryan Langholz, Argyrios Ziogas, Duncan C. Thomas, Cheryl Faucett, Mark Huberman, and Larry Goldstein. Ascertainment bias in rate ratio estimation from case-sibling control studies of variable age-at-onset diseases. Biometrics, 55(4):1129–1136, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gail:1999:TAE

[433] Mitchell H. Gail, Larry Kessler, Douglas Midthune, and Steven Scoppa. Two approaches for estimating disease prevalence from population-based registries of incidence and total mortality. Biometrics, 55(4):1137–1144, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Erkanli:1999:BIP

[434] Alaattin Erkanli, Refik Soyer, and Elizabeth J. Costello. Bayesian inference for prevalence in longitudinal two-phase studies. Biometrics, 55(4): 1145–1150, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:1999:MTI

[435] Dean A. Follmann and Michael A. Proschan. A multivariate test of interaction for use in clinical trials. Biometrics, 55(4):1151–1155, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shih:1999:CPT

[436] Joanna H. Shih and Michael P. Fay. A class of permutation tests for stratified survival data. Biometrics, 55(4):1156–1161, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carlin:1999:HPH

[437] Bradley P. Carlin and James S. Hodges. Hierarchical proportional hazards regression models for highly stratified data. Biometrics, 55(4):1162–1170, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Venkatraman:1999:PNT

[438] E. S. Venkatraman and Colin B. Begg. Properties of a nonparametric test for early comparison of treatments in clinical trials in the presence of surrogate endpoints. Biometrics, 55(4):1171–1176, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grigoletto:1999:ACI

[439] Matteo Grigoletto and Michael G. Akritas. Analysis of covariance with incomplete data via semiparametric model transformations. Biometrics, 55(4):1177–1187, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:1999:CTP

[440] Dei-In Tang and Nancy L. Geller. Closed testing procedures for group sequential clinical trials with multiple endpoints. Biometrics, 55(4):1188– 1192, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holcroft:1999:DVS

[441] Christina A. Holcroft and Donna Spiegelman. Design of validation studies for estimating the odds ratio of exposure-disease relationships when exposure is misclassified. Biometrics, 55(4):1193–1201, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chan:1999:TBE

[442] Ivan S. F. Chan and Zhongxin Zhang. Test-based exact confidence intervals for the difference of two binomial proportions. Biometrics, 55(4): 1202–1209, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bartoletti:1999:AE

[443] Stefania Bartoletti, Bernard D. Flury, and Daan G. Nel. Allometric extension. Biometrics, 55(4):1210–1214, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:1999:SIS

[444] Mei-Ling Ting Lee and G. A. Whitmore. Statistical inference for serial dilution assay data. Biometrics, 55(4):1215–1220, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mathew:1999:EMA

[445] Thomas Mathew and Kenneth Nordström. On the equivalence of metaanalysis using literature and using individual patient data. Biometrics, 55(4):1221–1223, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Satten:1999:VMB

[446] Glen A. Satten, Robert Janssen, Michael P. Busch, and Somnath Datta. Validating marker-based incidence estimates in repeatedly screened populations. Biometrics, 55(4):1224–1227, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Satten:1999:EET

[447] Glen A. Satten. Estimating the extent of tracking in interval-censored chain-of-events data. Biometrics, 55(4):1228–1231, December 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shih:1999:LMC

[448] Joanna H. Shih and Paul S. Albert. Latent model for correlated binary data with diagnostic error. Biometrics, 55(4):1232–1235, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:1999:NIM

[449] Yuh-Ing Chen. Nonparametric identification of the minimum effective dose. Biometrics, 55(4):1236–1240, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Regal:1999:ETA

[450] Ronald R. Regal and Ernest B. Hook. An exact test for all-way interaction in a 2^{M} contingency table: Application to interval capture–recapture estimation of population size. Biometrics, 55(4):1241–1246, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:1999:BRR

[451] Heping Zhang and Daniel Zelterman. Binary regression for risks in excess of subject-specific thresholds. Biometrics, 55(4):1247–1251, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:1999:MSM

[452] Paul S. Albert. A mover–stayer model for longitudinal marker data. Biometrics, 55(4):1252–1257, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:1999:RBT

[453] Yuh-Ing Chen. Rank-based tests for dose finding in nonmonotonic doseresponse settings. Biometrics, 55(4):1258–1262, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1999:EER

[454] You-Gan Wang. Estimating equations for removal data analysis. Biometrics, 55(4):1263–1268, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:1999:VDC

[455] John J. Chen and GlenysM Thomson. The variance for the disequilibrium coefficient in the individual Hardy-Weinberg test. Biometrics, 55(4):1269– 1272, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:1999:UHC

[456] Jianguo Sun. On the use of historical control data for trend test in carcinogenicity studies. Biometrics, 55(4):1273–1276, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lindsey:1999:MEC

[457] J. K. Lindsey. Multivariate elliptically contoured distributions for repeated measurements. Biometrics, 55(4):1277–1280, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Louzada-Neto:1999:PML

[458] Francisco Louzada-Neto. Polyhazard models for lifetime data. Biometrics, 55(4):1281–1285, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lehmacher:1999:ASS

[459] Walter Lehmacher and Gernot Wassmer. Adaptive sample size calculations in group sequential trials. Biometrics, 55(4):1286–1290, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kolassa:1999:SSC

[460] John E. Kolassa and Martin A. Tanner. Small-sample confidence regions in exponential families. Biometrics, 55(4):1291–1294, December 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wei:1999:MCI

[461] Wen Hsiang Wei and Ju Shiang Su. Model choice and influential cases for survival studies. Biometrics, 55(4):1295–1299, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fritsch:1999:MCE

[462] Kathleen S. Fritsch and Jason C. Hsu. Multiple comparison of entropies with application to dinosaur biodiversity. Biometrics, 55(4):1300–1305, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Auranen:1999:HBM

[463] Kari Auranen, Martin Eichner, Helena Käyhty, Aino K. Takala, and Elja Arjas. A hierarchical Bayesian model to predict the duration of immunity to Haemophilus influenzas Type B. Biometrics, 55(4):1306–1313, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Senn:1999:COA

[464] Stephen Senn and Andrew P. Grieve. A comment on optimal allocations for bioequivalence studies. Biometrics, 55(4):1314–1315, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1999:EB

[465] Anonymous. To the Editor of Biometrics. Biometrics, 55(4):1316–1317, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1999:ARF

[466] Anonymous. The authors replied as follows. Biometrics, 55(4):1316–1370, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kemp:1999:BRd

[467] C. D. Kemp and A. W. Kemp. Book reviews. Biometrics, 55(4):1318–1322, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1999:A

[468] Anonymous. Acknowledgements. Biometrics, 55(4):1323, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2003:CEI

[469] Anonymous. Correction to “Estimating Intraclass Correlation for Binary Data” by M. S. Ridout, C. G. B. Demétrio, and D. Firth; 55, 137–148, March 1999. Biometrics, 59(1):199, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [286].

Anonymous:2003:CEA

[470] Anonymous. Correction to “Estimation of absolute risk from nested casecontrol data” by B. Langholz and Ø. Borgan; 53, 767–774, June 1997. Biometrics, 59(2):451, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [221].

Anonymous:2013:CBM

[471] Anonymous. Correction to “Bayesian Monitoring of Event Rates with Censored Data,” by D. A. Follmann and P. S. Albert; 55, 603–607, DOI: 10.1111/j.0006-341X.1999.00603.x June 1999. Biometrics, 69(2):556, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [352].

Anonymous:2014:CUP

[472] Anonymous. Correction to “Using Pooled Exposure Assessment to Improve Efficiency in Case-Control Studies,” by Clarice R. Weinberg and David M. Umbach; 55, 718–726, September 1999. Biometrics, 70(4):1061, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [370].